### v4

# On the Schrödingerization method for linear non-unitary dynamics with optimal dependence on matrix queries

Abstract
The Schrödingerization method converts linear partial and ordinary differential equations with non-unitary dynamics into systems of Schrödinger-type equations with unitary evolution. It does so via the so-called warped phase transformation that maps the original equation into a Schrödinger-type equation in one higher dimension [Schrshort, JLY22SchrLong]. The original proposal used a particular initial function in the auxiliary space that did not achieve optimal scaling in precision. Here we show that, by choosing smoother initial functions in auxiliary space, Schrödingerization can in fact achieve near optimal and even optimal scaling in matrix queries. We construct three necessary criteria that the initial auxiliary state must satisfy to achieve optimality. This paper presents detailed implementation of four smooth initializations for the Schrödingerization method: (a) the error function and related functions, (b) the cut-off function, (c) the higher-order polynomial interpolation, and (d) Fourier transform methods. Method (a) achieves optimality and methods (b), (c) and (d) can achieve near-optimality. A detailed analysis of key parameters affecting time complexity is conducted.
1 Introduction
Quantum computing, an emerging technology, utilizes the principles of quantum mechanics to achieve unprecedented computational power [hidary2019quantum, NC2010Quantum, preskill2018quantum, rieffel2011quantum]. Quantum algorithms operate within an $n$-qubit Hilbert space of dimension $2^{n}$, potentially offering polynomial to even exponential computational advantage for models involving vast amounts of data. Hence, it has become an attractive computational paradigm to handle large-scale scientific computing problems that are bottlenecks for classical computation. A natural application is partial differential equations (PDEs) from time-dependent Schrödinger equations, which follow unitary evolutions and hence the wave functions can be coherently represented on quantum computers. Known as Hamiltonian simulations, a variety of efficient algorithms have been developed toward this goal [BCC15, Berry-Childs-Kothari-2015, BCC17, LC17, Low2019qubitization, CGJ19, Berry2019Dyson, AnLin2021TimeHamiltonian, JLY22SchrLong, An2022timedependent, fang2023time].
However, many physical phenomena – such as combustion, atmospheric and oceanic circulation, and electromagnetic wave propagation with physical boundaries – exhibit non-unitary dynamics. Even the time-dependent Schrödinger equation becomes non-unitary when artificial boundary conditions are introduced, making traditional Hamiltonian simulation techniques inapplicable [JLLY23ABC]. Alternatively, crafting quantum PDE solvers involves discretizing spatial variables to formulate a system of ordinary differential equations (ODEs), which can then be tackled using quantum ODE solvers [Berry-2014, BerryChilds2017ODE, Childs-Liu-2020]. In recent years, significant progress has been made in designing and analyzing efficient quantum algorithms for linear ODEs. These algorithms can be classified into several categories. The first category involves three steps: discretizing the time variable, encoding the discretized linear differential equation into an enlarged linear system, and solving the resulting system using Quantum Linear System Algorithms (QLSA) [Berry-2014, BerryChilds2017ODE, KroviODE, JLY2022multiscale, WL24, DLX25, Childs-Liu-2020, Berry2024Dyson]. The second category leverages an integral representation of the non-unitary evolution operator, followed by Linear Combination of Hamiltonian Simulations [ALL2023LCH, ACL2023LCH2], and its relationship to Schrödingerization, the method in the third category, is explored in [Li25]. The third category involves dilating the system into a unitary system. Crucially, if the solution operator for the resulting ODE system is unitary, quantum simulations can achieve reduced time complexity compared to other quantum linear algebra methods [Berry-2014, JLY2022multiscale, ALWZ25, WL24, DLX25]. In cases where the system is not unitary – such as when incorporating physical boundary conditions – it becomes necessary to “dilate” it to a unitary system [JLY22SchrLong, Javier2022optionprice, Burgarth23dilations, CJL23TimeSchr, DLX25]. Other methods, such as those employing block-encoding techniques, can be found in [ALWZ25, fang2023time].
Among the unitarization techniques, the Schrödingerization method proposed in [Schrshort, JLY22SchrLong] offers a simple and general framework enabling quantum simulation for all linear PDEs and ODEs. It employs a warped phase transformation to lift the original equations into a higher-dimensional space, where they become Schrödinger-type equations – with unitary evolutions – in the Fourier domain! This method has since been extended to a wide range of settings, including open quantum systems with non-unitary artificial boundary conditions [JLLY23ABC], systems with physical boundary and interface conditions [JLLY2024boundary], Maxwell’s equations [JLC23Maxwell, MJLWZ24], the Fokker-Planck equation [JLY24FokkerPlanck], ill-posed problems such as the backward heat equation [JLM24SchrBackward]. It has also been applied in iterative linear algebra solvers [JinLiu-LA]. Moreover, as a naturally continuous-variable method [analogPDE], it represents the only viable approach so far for analog quantum simulation of PDEs and ODEs. The method can also be adapted to parabolic PDEs using a Jaynes-Cummings-type model, which is more readily available on current devices  [JL24JaynesCummings].
We consider linear dynamical systems with a general evolution operator $A(t)$ and an inhomogeneous term $\bm{b}(t)$ as given in (2.1), and present a detailed implementation of the corresponding quantum algorithm using block-encoding techniques as described in [Low2019Interaction, ACL2023LCH2]. In our implementation, we first transform the system into a homogeneous one by enlarging the system with an auxiliary variable. Using the Schrödingerization method, we then transform it into a Hamiltonian system with unitary evolution operator, which can be efficiently solved on quantum computers. For time-dependent Hamiltonian dynamics, we apply the quantum simulation technique from [CJL23TimeSchr, CJL24] for non-autonomous systems. This technique involves transforming a non-autonomous system into an autonomous one in a higher dimension, avoiding the need for the complicated Dyson series. Consequently, we focus exclusively on analyzing the optimal dependence in the time-independent case.
In this article, we focus on the optimal scaling behavior of the Schrödingerization method for non-unitary dynamics. In the original Schrödingerization method, a simple even-function $\psi(p)=\mathrm{e}^{-|p|}$ is used as the initial auxiliary state, resulting only in a first-order approximation, due to the lack of regularity of this state. This lack of regularity meant that achieving precision $\epsilon$ may require a small enough mesh size $\triangle p=\mathcal{O}(\epsilon)$, which results in the maximum absolute value among the discrete Fourier modes scaling being $\mathcal{O}(1/\epsilon)$, i.e., $\mu_{\max}=\mathcal{O}(1/\epsilon)$. This is not optimal because the query complexity linearly depends on $\mu_{\max}$, which is the maximum Fourier mode in absolute value, for the extended variable $p$.
Since this non-optimal $\mathcal{O}(1/\epsilon)$ scaling is ultimately due to the lack of regularity of the initial auxiliary state function, then improving this scaling is only a question of how to smooth out this initial function in an appropriate way. For example, as already pointed out in [JLM24SchrBackward], we can achieve improved – near optimal – scaling by employing a smoother initialization. In this article we provide four different smooth initializations, with detailed analysis on their complexities. One of these methods (Section 4) leads to optimal complexity and the other three methods (Section 3) lead to near optimal complexity. In Section 4 we also provide three necessary criteria that $\psi(p)$ should satisfy in order to achieve optimal scaling.
Through a detailed analysis of the parameters affecting time complexity, we find that the query complexity scales linearly with $\mu_{\max}$, the maximum Fourier mode in absolute value, for the extended variable $p$ with a generic initialization function $\psi(p)$ in the $p$-domain. Assuming that $\|\psi-P_{h}\psi\|_{L^{2}(\mathbb{R})}\leq\epsilon$ for $\psi(p)\in H^{r}(\mathbb{R})$, with $P_{h}\psi$ being the discrete Fourier approximation of $\psi$, we observe that $\mu_{\max}$ scales as $\mathcal{O}((1/\epsilon)^{1/r}\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r})$, where $\psi^{(r)}$ denotes the $r$-th derivative of $\psi$ and $H^{r}$ is the standard Sobolev space. When $r$ is sufficiently large, specifically $r\simeq\log(1/\epsilon)$, we have $\mu_{\max}\simeq\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}$. Thus, the precision scaling problem reduces to identifying an appropriate function $\psi(p)$ such that
$\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\leq Cr^{1/\beta},\qquad\beta\in(0,1],$ (1.1)
where $\beta=1$ implies the optimal precision dependence.
Based on this observation, we provide an abstract framework for the complexity analysis in Theorem 2.2. For academic interest, we first provide several sufficiently smooth initializations that offer nearly exponential speedup in the $p$-variable over the original Schrödingerization method in terms of precision $\varepsilon$, including cut-off functions, higher-order interpolation, and Fourier transform methods. Since we impose the condition $\psi(p)=\mathrm{e}^{-p}$ for $p>0$, the dependence on matrix queries can only be made near-optimal, as $\beta<1$. Achieving optimal dependence requires $\beta=1$. We therefore summarize the conditions on $\psi$ necessary for attaining optimal dependence and provide a detailed estimate. In particular, we employ the error function $\text{erf}(p)$ to achieve the required optimal bound and show that this is not the unique function, but is the simplest.
Our construction is based on a simple yet important observation: the process of constructing a smooth initialization function $\psi(p)$ reduces to finding a smooth approximation of the step function. Such smoothness is typically achieved through convolution. Given the periodic boundary conditions required for the discrete Fourier transform, we first consider using a mollifier as the convolution kernel. The mollifier is infinitely smooth and has compact support, thereby realizing the exact periodic boundaries. However, upon careful analysis, we find that smoothness alone is insufficient to satisfy (4.3). In fact, we require a stricter condition: the step function approximation must also be analytic. Under this condition, the most natural choice for the convolution kernel is the Gaussian $\mathrm{e}^{-p^{2}}$, which leads to the expression involving the error function. For both choices of convolution kernels, we provide a rigorous proof of the bounds for the derivatives of the smooth initialization functions. This demonstrates that the mollifier yields $\beta=1/2$, while the Gaussian results in $\beta=1$. Consequently, we derive the near-optimal cost for the mollifier and the optimal scaling behavior for the Gaussian by applying the abstract framework for complexity analysis.
In Table 1, we compare our algorithm with previous approaches in the homogeneous case. It is evident that the Schrödingerization method, with sufficiently smooth initialization in $p$, achieves both optimal state preparation cost and optimal dependence of the number of queries to the matrix on all parameters. We also note that the improvement in the LCHS method in [ACL2023LCH2] only leads to a sub-optimal dependence on matrix queries, as $\beta$ cannot be exactly equal to 1 (i.e., $1/\beta>1$), where $\beta$ is the parameter in the kernel function of the improved LCHS.
During the revision stage of this work, we became aware of a concurrent study by [Low25LCHS], which presents a generalized approximate version of LCHS. This extension incorporates a kernel function with exponential decay, enabling a quantum ODE algorithm that achieves optimal dependence on precision.
Table 1: Comparison among improved Schrödingerization and previous methods for homogeneous dynamical systems $\mathrm{d}\bm{u}/\mathrm{d}t=A\bm{u}$. Here, $u_{r}=\frac{\|\bm{u}_{0}\|}{\|\bm{u}(T)\|}$, $\alpha_{A}\geq\|A\|$, $T$ is the evolution time, $\varepsilon$ is the error, and $\beta\in(0,1)$. All but the spectral method assume the real part of AA to be negative semi-definite, while in the spectral method AA is assumed to be diagonalizable with matrix $V$ such that $\kappa_{V}\geq\|V^{-1}\|\|V\|$ and all the eigenvalues of AA have non-positive real parts.
Method Queries to AA Queries to $\bm{u}_{0}$
Spectral method [Childs-Liu-2020] $\widetilde{\mathcal{O}}(u_{r}\kappa_{V}\alpha_{A}T\text{poly}(\log\frac{1}{\varepsilon}))$ $\widetilde{\mathcal{O}}(u_{r}\kappa_{V}\alpha_{A}T\text{poly}(\log\frac{1}{\varepsilon}))$
Truncated Dyson [Berry2019Dyson] $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T(\log\frac{1}{\varepsilon})^{2})$ ${\mathcal{O}}(u_{r}\alpha_{A}T\log\frac{1}{\varepsilon})$
Time-marching [fang2023time] $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}^{2}T^{2}\log\frac{1}{\varepsilon})$ ${\mathcal{O}}(u_{r})$
Improved LCHS time-dependent [ACL2023LCH2] $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T(\log\frac{1}{\varepsilon})^{1+1/\beta})$ ${\mathcal{O}}(u_{r})$
Improved LCHS time-independent [ACL2023LCH2] $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T(\log\frac{1}{\varepsilon})^{1/\beta})$ ${\mathcal{O}}(u_{r})$
Optimal LCHS time-independent [Low25LCHS] $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T\log\frac{1}{\varepsilon})$ ${\mathcal{O}}(u_{r})$
This work, time-dependent $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T(\log\frac{1}{\varepsilon})^{2})$ ${\mathcal{O}}(u_{r})$
This work, time-independent $\widetilde{\mathcal{O}}(u_{r}\alpha_{A}T\log\frac{1}{\varepsilon})$ ${\mathcal{O}}(u_{r})$
Notation.
Throughout the paper, we adopt zero-based indexing: indices $j\in\{0,1,\dots,N-1\},$ and we also write $j\in[N]$ with $N]:=\{0,1,\dots,N-1\}$. We use $|j\rangle\in\mathbb{C}^{N}$ to denote the $j$-th canonical basis vector, whose $j$-th component is $1$ and all other components are 0. We denote the identity and zero matrices by $I$ and $O$, respectively; their dimensions should be clear from the context. When clarification is needed, we write $I_{N}$ for the $N\times N$ identity matrix. In particular, $\mathbf{1}$ denotes the $2\times 2$ identity matrix (acting on a single qubit).
Vector-valued quantities are denoted by boldface symbols, e.g. $\bm{u}$. Given a nonzero vector $\bm{u}$, the notation $|\bm{u}\rangle$ represents the pure quantum state obtained by normalizing $\bm{u}$ in the Euclidean norm, $|\bm{u}\rangle\;=\;\frac{\bm{u}}{\|\bm{u}\|}.$
Unless otherwise specified, for a vector $\bm{u}\in\mathbb{C}^{N}$ we write $\|\bm{u}\|:=\|\bm{u}\|_{2}$ for the standard Euclidean ($l^{2}$) norm. For a matrix $A\in\mathbb{C}^{N\times N}$, the notation $\|A\|$ refers to the operator norm induced by the Euclidean norm, $\|A\|:=\sup_{\bm{u}\neq 0}\frac{\|A\bm{u}\|}{\|\bm{u}\|}.$ When Sobolev norms are used, we write $\|\cdot\|_{H^{s}(\Omega)}$, $\|\cdot\|_{L^{2}(\Omega)}$, etc., and these always denote the standard norms on the corresponding Sobolev spaces.
[…]
Organization of the paper.
The paper is structured as follows. In Section 2, we offer an overview of the Schrödingerization approach, present the full implementation by using block-encoding techniques and establish an abstract framework for the complexity analysis. Section 3 provide sufficiently smooth initializations that offer nearly exponential speedup in the $p$-variable over the original Schrödingerization method. Section 4 demonstrates how optimal scaling in matrix queries can be attained through the modification of the initializations of the warped phase transformation. We establish the optimal dependence by constructing a function using the error function $\text{erf}(p)$. Section 5 shows the detailed error estimate for the Schrödingerization. Finally, some discussions are presented in the last section.
2 The Schrödingerization method for non-unitary dynamics
Consider a system of linear dynamical system in the form
$\begin{cases}\dfrac{\mathrm{d}}{\mathrm{d}t}\bm{u}(t)=A(t)\bm{u}(t)+\bm{b}(t),\qquad t\in(0,T),\\ \bm{u}(0)=\bm{u}_{0},\end{cases}$ (2.1)
where $T$ is the evolution time, $\bm{u}=[u_{0},u_{1},\cdots,u_{N-1}]^{\top},\bm{b}=[b_{0},b_{1},\cdots,b_{N-1}]^{\top}\in\mathbb{C}^{N}$ and $A\in\mathbb{C}^{N\times N}$. In general, $A$ is not anti-Hermitian, i.e., $A^{\dagger}\neq-A$, where $\dagger$” denotes conjugate transpose. When $A$ is a linear operator, (2.1) is a system of ODEs. When $A$ is a linear differential operator, (2.1) is a system of PDEs. By introducing an auxiliary vector function $\bm{r}(t)$ that remains constant in time if $\bm{b}\neq 0$, system (2.1) can be rewritten as a homogeneous system
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{u}_{f}=A_{f}\bm{u}_{f},\quad A_{f}=\begin{bmatrix}A&B\\ O&O\end{bmatrix},\quad\bm{u}_{f}(0)=\bm{u}_{I}:=\begin{bmatrix}\bm{u}_{0}\\ \bm{r}_{0}\end{bmatrix}$, (2.2)
where $B=\text{diag}\{b_{0}/\gamma_{0},\cdots,b_{N-1}/\gamma_{N-1}\}$ and $\bm{r}_{0}=[\gamma_{0},\cdots,\gamma_{N-1}]^{\top}$, with
$\gamma_{i}=T\sup_{t\in[0,T]}|b_{i}(t)|,\qquad i=0,1,\cdots,N-1.$ (2.3)
Here, each $\sup_{t\in[0,T]}|b_{i}(t)|$ can be replaced by its upper bound and we set $b_{i}/\gamma_{i}=0$ if $b_{i}(t)\equiv 0$.
2.1 The Schrödingerizaton method
In this section, we briefly review the Schrödingerization approach for general linear dynamical systems. For a general $A_{f}$, we first decompose $A_{f}$ into a Hermitian term and an anti-Hermitian term:
$A_{f}(t)=H_{1}(t)+\mathrm{i}H_{2}(t),\qquad\mathrm{i}=\sqrt{-1},$
where
$H_{1}(t)=\frac{A_{f}(t)+A_{f}^{\dagger}(t)}{2}=\begin{bmatrix}H_{1}^{A}&\frac{B}{2}\\ \frac{B^{\top}}{2}&O\end{bmatrix},\quad H_{2}(t)=\frac{A_{f}(t)-A_{f}^{\dagger}(t)}{2\mathrm{i}}=\begin{bmatrix}H_{2}^{A}&\frac{B}{2\mathrm{i}}\\ -\frac{B^{\top}}{2\mathrm{i}}&O\end{bmatrix}$,
with $H_{1}^{A}=(A+A^{\dagger})/2$ and $H_{2}^{A}=(A-A^{\dagger})/(2\mathrm{i})$. Throughout the article, we assume that the real part matrix $H_{1}^{A}$ is negative semi-definite. More general cases are addressed in [JLM24SchrInhom, JLM24SchrBackward, JLY24FokkerPlanck].
Using the warped phase transformation $\bm{w}(t,p)=\mathrm{e}^{-p}\bm{u}_{f}(t)$ for $p\geq 0$ and symmetrically extending the initial data to $p<0$, system (2.1) is then transformed to a system of linear convection equations [Schrshort, JLY22SchrLong]:
$\begin{cases}\dfrac{\partial}{\partial t}\bm{w}(t,p)=-H_{1}(t)\partial_{p}\bm{w}+\mathrm{i}H_{2}(t)\bm{w},\\ \bm{w}(0,p)=\psi(p)\bm{u}_{I},\end{cases}$ (2.4)
where $\psi(p):=\mathrm{e}^{-|p|}$. According to [JLM24SchrInhom, Theorem 3.1], we can restore the solution $\bm{u}_{f}(t)$ by
$\bm{u}_{f}=\mathrm{e}^{p}\bm{w}(t,p),\quad p\geq p^{\Diamond}=\lambda_{\max}^{+}(H_{1})T.$ (2.5)
Here $\lambda_{\max}^{+}(H_{1})$ is defined by
$\lambda_{\max}^{+}(H_{1})=\begin{cases}\displaystyle\sup_{\begin{subarray}{c}\lambda\in\sigma(H_{1}(t))\ 0<t<T,0<\lambda\end{subarray}}|\lambda|,&\text{if }\exists\lambda>0\text{ in }\sigma(H_{1}(t))\text{ over }[0,T),\\ 0,&\text{otherwise},\end{cases}$ (2.6)
with $\sigma(H_{1})$ the set of eigenvalues of $H_{1}$. Similarly, $\lambda_{\max}^{-}(H_{1})$ is defined by
$\lambda_{\max}^{-}(H_{1})=\begin{cases}\displaystyle\sup_{\begin{subarray}{c}\lambda\in\sigma(H_{1}(t))\ 0<t<T,\lambda<0\end{subarray}}|\lambda|,&\text{if }\exists\lambda<0\text{ in }\sigma(H_{1}(t))\text{ over }[0,T),\\ 0,&\text{otherwise}.\end{cases}$ (2.7)
Since $H_{1}^{A}$ is negative, it is easy to find from (2.3) that
$\lambda_{\max}^{+}(H_{1})T\leq\frac{1}{2}\|B\|_{\max}T\leq\frac{1}{2}.$
For numerical implementation, we truncate the extended region to a finite interval $p\in[-L,R]$ with $L>0$ and $R>0$ satisfying
$\mathrm{e}^{-L+\lambda_{\max}(H_{1})T}\approx\mathrm{e}^{-R+\lambda_{\max}(H_{1})T}\approx\epsilon.$ (2.8)
Here $\lambda_{\max}(H_{1})$ denotes the largest absolute value among the eigenvalues of $H_{1}$, and $\epsilon$ is a predetermined tolerance constant, which will be specified later.
The requirement in (2.8) is explained as follows. Since the original problem is posed on the whole space, we truncate it to a finite interval [L,R][-L,R] with periodic boundary conditions. This means that we require $\bm{w}(0,-L)\approx\bm{w}(0,R)\approx\epsilon$, or equivalently, $\mathrm{e}^{-L}\approx\mathrm{e}^{-R}\approx\epsilon$. For the transport equation $u_{t}-au_{p}=0$ with $a>0$, the initial value at $p_{0}$, i.e., $u(0,p_{0})$, will remain constant along the characteristic line $p+at=p_{0}$, which implies $u(t,p_{0}-at)=u(0,p_{0})$. For the transport equation in (2.4) with periodic boundaries, the solution values at p=Lp=-L and $p=R$ must also be compatible along characteristics, based on the initial data in the regions $(-L,-L+\lambda_{\max}(H_{1})T)$ and $(R-\lambda_{\max}(H_{1})T,R)$. If the initial data in these regions has already decayed to the level of $\epsilon$, then the boundary values and their higher derivatives satisfy
$\bm{w}^{(k)}(t,-L)\approx\bm{w}^{(k)}(t,R)\approx\epsilon,\quad t\in[0,T],$
so that the periodic boundary condition is consistent with the infinite-domain problem up to accuracy $\epsilon$.
Toward this end, we choose a uniform mesh size $\triangle p=(R+L)/N_{p}$ for the auxiliary variable with $N_{p}=2^{n_{p}}$ being an even number, with the grid points denoted by $L=p_{0}<p_{1}<\cdots<p_{N_{p}}=R$. Let the vector $\bm{W}_{h}\in\mathbb{C}^{N_{np}\times 1}$ with $N_{np}=N\times N_{p}$ be the collection of $\bm{w}_{h}(t,p)$ at these grid points, defined more precisely as $\bm{W}_{h}(t)=\sum_{k\in[N_{p}],i\in[N]}w_{i,h}(t,p_{k})|k,i\rangle$, where $w_{i,h}$ is the $i$-th entry of $\bm{w}_{h}$ and $|k,i\rangle=|k\rangle|i\rangle$.
By applying the discrete Fourier transform in the $p$ direction, one arrives at
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{W}_{h}(t)=-\mathrm{i}(P_{\mu}\otimes H_{1})\bm{W}_{h}+\mathrm{i}(I\otimes H_{2})\bm{W}_{h},\quad\bm{W}_{h}(0)=\bm{\psi}\otimes\bm{u}_{I},$ (2.9)
where $\bm{\psi}=[\psi(p_{0}),\cdots,\psi(p_{N_{p}-1})]^{\top}$. Here, $P_{\mu}$ is the matrix expression of the momentum operator $\mathrm{i}\partial_{p}$, given by
$P_{\mu}=\Phi D_{\mu}\Phi^{-1},\qquad D_{\mu}=\text{diag}(\mu_{0},\cdots,\mu_{N_{p}-1}),$ (2.10)
where $\mu_{k}=\frac{2\pi}{R+L}(k-\frac{N_{p}}{2})$ are the Fourier modes and
$\Phi=(\phi_{jl})_{N_{p}\times N_{p}}=(\phi_{l}(x_{j}))_{N_{p}\times N_{p}},\qquad\phi_{l}(x)=\mathrm{e}^{\mathrm{i}\mu_{l}(x+L)}$
is the matrix representation of the discrete Fourier transform. At this point, we have successfully mapped the dynamics back to a Hamiltonian system. By a change of variables $\tilde{\bm{W}}_{h}=(\Phi^{-1}\otimes I)\bm{W}_{h}$, one has
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{\bm{W}}_{h}(t)=-\mathrm{i}H(t)\tilde{\bm{W}}_{h}(t),$ (2.11)
where $H=D_{\mu}\otimes H_{1}-I\otimes H_{2}$.
Remark 2.1.
Our method for solving (2.1) with a time-dependent source term encodes $\bm{b}(t)$ directly within the coefficient matrix. This results in the same query complexity for both the coefficient matrix $A(t)$ and $\bm{b}(t)$. To minimize the repeated use of the state preparation oracle $O_{b}$ for the source term $\bm{b}$, when $\bm{b}(t)$ is time-independent, we can instead consider a simpler enlarged system
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{u}_{f}(t)=\begin{bmatrix}A&\frac{I}{T}\\ O&O\end{bmatrix}\bm{u}_{f}(t),\qquad\bm{u}_{f}(t)=\begin{bmatrix}\bm{u}(t)\\ T\bm{b}\end{bmatrix},\qquad\bm{u}_{f}(0)=\bm{u}_{I}:=\begin{bmatrix}\bm{u}(0)\\ T\bm{b}\end{bmatrix}$.
In the time-dependent case, it may be advantageous to separately implement their homogeneous and inhomogeneous parts and combine them using the LCU technique [ACL2023LCH2]. Each execution of the LCU procedure requires $\mathcal{O}(1)$ uses of the associated preparation oracles, with the overall complexity primarily dependent on the success probability.
2.2 Quantum simulation for time-dependent Schrödingerized system
If the coefficient in the dynamical system is time-dependent, namely a non-automomous system, one can turn it into an autonomous unitary system via dimension lifting [CJL23TimeSchr]. First, via Schrödingerization, one obtains a time-dependent Hamiltonian
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{\bm{W}}_{h}=-\mathrm{i}H(t)\tilde{\bm{W}}_{h},\quad H=H^{\dagger}.$ (2.12)
By introducing a new “time” variable $s$, the problem becomes a new linear PDE defined in one higher dimension but with time-independent coefficients,
$\frac{\partial\bm{v}}{\partial t}=-\frac{\partial\bm{v}}{\partial s}-\mathrm{i}H(s)\bm{v}\qquad\bm{v}(0,s)=\delta(s)\tilde{\bm{w}}_{h}(0),\quad s\in\mathbb{R},$ (2.13)
where $\delta(s)$ is the dirac $\delta$-function. One can easily recover $\tilde{\bm{W}}_{h}$ by $\tilde{\bm{W}}_{h}=\int_{-\infty}^{\infty}\bm{v}(t,s)\;\mathrm{d}s$.
Since $\bm{v}$ decays to zero as $s$ approaches infinity, the $s$-region can be truncated to $\pi S,\pi S]$, where $\pi S>4\omega+T$, with $2\omega$ representing the length of the support set of the approximated delta function. Choosing $S$ sufficiently large ensures that the compact support of the approximated delta function remains entirely within the computational domain throughout the simulation, allowing the spectral method to be applied. The transformation and difference matrix are defined by
$(\Phi_{s})_{lj}=(\mathrm{e}^{\mathrm{i}\mu^{s}_{l}(j\triangle s)}),\quad D_{s}=\text{diag}\{\mu^{s}_{0},\mu^{s}_{1},\dots,\mu^{s}_{N_{s}-1}\},\quad\mu^{s}_{l}=(l-\frac{N_{s}}{2})S,\quad l,j\in[N_{s}],$
where $\triangle s=2\pi S/N_{s}$. Applying the discrete Fourier spectral discretization, it yields a time-independent Hamiltonian system as
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{\bm{V}}_{h}=-\mathrm{i}\big(D_{s}\otimes I+I_{N_{s}}\otimes H\big)\tilde{\bm{V}}_{h},\quad\tilde{\bm{V}}_{h}(0)=[\Phi_{s}^{-1}\otimes I](\mathbf{\delta}_{\bm{h}}\otimes\tilde{\bm{W}}_{h}(0)),$ (2.14)
where $\mathbf{\delta}_{\bm{h}}=\sum\limits_{j\in[N_{s}]}\delta_{w}(s_{j})|j\rangle$ with $s_{j}=-\pi S+j\triangle s$ and $\delta_{\omega}$ is an approximation to $\delta$ function defined, for example, by choosing
$\delta_{\omega}(x)=\frac{1}{\omega}\bigg(1-\frac{1}{2}|1+\cos(\pi\frac{x}{\omega})|\bigg)\quad|x|\leq\omega,\quad\delta_{\omega}(x)=0\quad|x|\geq\omega.$
Here $\omega=m\triangle s$, where $m$ is the number of mesh points within the support of $\delta_{\omega}$.
2.3 Hamiltonian system for quantum computing
As discussed in Section 2.2, a time-dependent system can be transformed into a time-independent system by adding an additional dimension. Therefore, the subsequent discussion will focus exclusively on the time-independent case. For further details on time-dependent systems, we refer to [CJL24].
From (2.11), a quantum simulation can be carried out on the Hamiltonian system above:
$|\bm{W}_{h}(T)\rangle=\big[\Phi\otimes I\big]\cdot\mathcal{U}(T)\cdot\big[\Phi^{-1}\otimes I\big]|\bm{W}_{h}(0)\rangle,$
where $\mathcal{U}(T)=\mathrm{e}^{-\mathrm{i}HT}$ is a unitary operator, and $\Phi$ (or $\Phi^{-1}$) is completed by (inverse) quantum Fourier transform (QFT or IQFT). The complete circuit for implementing the quantum simulation of $|\bm{w}_{h}\rangle$ is illustrated in Fig. 1.
Fig. 1: Quantum circuit for Schrödingerization of  (2.11), where $\bm{\psi}_{h}=\sum_{k\in[N_{p}]}\psi(p_{k})|k\rangle$.
From (2.5), one can recover the target variables for $\bm{u}_{f}$ by performing a measurement in the computational basis:
$M_{k}=|k\rangle\langle k|\otimes I,\quad k\in\{j:p_{j}\geq p^{\Diamond}\;\text{and}\;p_{j}=\mathcal{O}(1)\}=:I_{\Diamond},$
where $I_{\Diamond}$ is referred to as the recovery index set. The state vector is then collapsed to
$|\bm{w}_{*}\rangle\equiv|k_{*}\rangle\otimes\frac{1}{\mathcal{N}}\Big(\sum\limits_{i}w_{k_{*}i}|i\rangle\Big),\quad\mathcal{N}=\Big(\sum\limits_{i}|w_{k_{*}i}|^{2}\Big)^{1/2},$
where $w_{k_{*},i}=\langle k_{*}|\langle i|\otimes\bm{W}_{h}$ for some $k_{*}$ in the recovery index set $I_{\Diamond}$ with the probability
$\text{P}_{\text{r}}(\bm{w}(T,p_{k_{*}}))=\frac{\sum_{i}|w_{k_{*}i}(T)|^{2}}{\sum_{k,i}|w_{ki}(T)|^{2}}=\frac{\|\bm{w_{*}}(T)\|^{2}}{\|\bm{w}_{h}(T)\|^{2}}=\frac{\|\bm{w_{*}}(T)\|^{2}}{\|\bm{w}_{h}(0)\|^{2}}.$
Then the likelihood of acquiring $|\bm{w}_{*}\rangle$ that satisfies $k_{*}\in I_{\Diamond}$ is given by
$\text{P}_{\text{r}}(\bm{w})=\sum_{k\in I_{\Diamond}}\text{P}_{\text{r}}(\bm{w}(T,p_{k}))=\frac{\sum_{k\in I_{\Diamond}}\sum_{i\in[N]}|w_{k,i}(T)|^{2}}{\|\bm{w}_{h}(0)\|^{2}}=\frac{C_{e0}^{2}}{C_{e}^{2}}\frac{\|\bm{u}_{f}(T)\|^{2}}{\|\bm{u}_{I}\|^{2}},$ (2.15)
where
$C_{e0}=\Big(\sum_{p_{k}\in I_{\Diamond}}(\psi(p_{k}))^{2}\Big)^{1/2},\qquad C_{e}=\Big(\sum_{k=0}^{N_{p}-1}(\psi(p_{k}))^{2}\Big)^{1/2},\qquad\psi(p)=\mathrm{e}^{-|p|}.$ (2.16)
If $N_{p}$ is sufficiently large, we have
$\triangle pC_{e0}^{2}\approx\int_{p^{\Diamond}}^{\infty}\mathrm{e}^{-2p}\mathrm{d}p=\frac{1}{2}\mathrm{e}^{-2p^{\Diamond}},\qquad\triangle pC_{e}^{2}\approx\int_{-\infty}^{\infty}\mathrm{e}^{-2p}\mathrm{d}p=1,$
where $\triangle p=(R+L)/N_{p}$, then it yields
$\frac{C_{e0}^{2}}{C_{e}^{2}}\approx\frac{1}{2}\mathrm{e}^{-2p^{\Diamond}}\geq\frac{1}{2}\mathrm{e}^{-1}.$ (2.17)
Since $\bm{u}_{f}(t)=\begin{bmatrix}\bm{u}(t)\\ \bm{r}_{0}\end{bmatrix}$, one can perform a projection to get $|\bm{u}(T)\rangle$ with the probability $\dfrac{\|\bm{u}(T)\|^{2}}{\|\bm{u}_{f}(T)\|^{2}}$. The overall probability of retrieving $\bm{u}$ is then approximated by
$\text{P}_{\text{r}}(\bm{u})=\text{P}_{\text{r}}(\bm{w})\cdot\frac{\|\bm{u}(T)\|^{2}}{\|\bm{u}_{f}(T)\|^{2}}=\frac{C_{e0}^{2}}{C_{e}^{2}}\frac{\|\bm{u}(T)\|^{2}}{\|\bm{u}_{I}\|^{2}}=\frac{C_{e0}^{2}}{C_{e}^{2}}\frac{\|\bm{u}(T)\|^{2}}{\|\bm{u}_{0}\|^{2}+T^{2}\|\bm{b}\|_{\mathrm{smax}}^{2}},$
where
$\|\bm{b}\|_{\mathrm{smax}}^{2}=\sum_{i=0}^{N-1}\Big(\sup_{t\in[0,T]}|b_{i}(t)|\Big)^{2}.$ (2.18)
By using the amplitude amplification, the repeated times for the measurements can be approximated as
$g=\mathcal{O}\Big(\frac{C_{e}}{C_{e0}}\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}\Big)=\mathcal{O}\Big(\frac{\|\bm{u}_{0}\|+T\|\bm{b}\|_{\mathrm{smax}}}{\|\bm{u}(T)\|}\Big).$ (2.19)
The quantity $g$ in (2.19) is comparable to the number of repeated times by directly projecting onto $|k_{*}\rangle|0\rangle$ for $k_{*}\in I_{\Diamond}$.
2.4 Detailed implementation of the Hamiltonian simulation
For the Hamiltonian simulation of $\mathcal{U}(T)=\mathrm{e}^{-\mathrm{i}HT}$, where $H$ arises from the time-dependent system, one can apply established quantum algorithms from the literature. For instance, Hamiltonian simulation with nearly optimal dependence on all parameters is discussed in [Berry-Childs-Kothari-2015], where sparse access to the Hamiltonian $H$ is assumed.
One can express the evolution operator $\mathcal{U}(T)$ as a select oracle
$\mathcal{U}(T)=\sum_{k=0}^{N_{p}-1}|k\rangle\langle k|\otimes\mathrm{e}^{-\mathrm{i}(\mu_{k}H_{1}-H_{2})T}=:\sum_{k=0}^{N_{p}-1}|k\rangle\langle k|\otimes V_{k}(T).$
Since the unitary $V_{k}(T)$ corresponds to the simulation of the Hamiltonian $H_{\mu_{k}}:=\mu_{k}H_{1}-H_{2}$, we assume the block-encoding oracles encoding the real and imaginary parts separately, namely
$(\langle 0|_{a}\otimes I)U_{H_{i}}(|0\rangle_{a}\otimes I)=\frac{H_{i}}{\alpha_{i}},\qquad i=1,2,$
where $\alpha_{i}\geq\|H_{i}\|$ is the block-encoding factor for $i=1,2$.
According to the discussion in [ACL2023LCH2, Section 4.2.1], there is an oracle $\text{HAM-T}_{H_{\mu}}$ such that
$(\langle 0|_{a^{\prime}}\otimes I)\text{HAM-T}_{H_{\mu}}(|0\rangle_{a^{\prime}}\otimes I)=\sum_{k=0}^{N_{p}-1}|k\rangle\langle k|\otimes\frac{H_{\mu_{k}}}{\alpha_{1}\mu_{\max}+\alpha_{2}},$ (2.20)
where $H_{\mu_{k}}=\mu_{k}H_{1}-H_{2}$ and $\mu_{\max}=\max_{k}|\mu_{k}|$ represent the maximum absolute value among the discrete Fourier modes. This oracle only uses $\mathcal{O}(1)$ queries to block-encoding oracles for $H_{1}$ and $H_{2}$. With the block-encoding oracle $\text{HAM-T}_{H_{\mu}}$, we can implement
$\text{SEL}_{0}=\sum_{k=0}^{N_{p}-1}|k\rangle\langle k|\otimes V_{k}^{a}(T),$
a block-encoding of $\mathcal{U}(T)$, using the quantum singular value transformation (QSVT) [gilyen2019quantum] for example, where $V_{k}^{a}(T)$ block-encodes $V_{k}(T)$ with
$\|V_{k}^{a}(T)-V_{k}(T)\|\leq\delta.$ (2.21)
This uses the oracles for $H_{1}$ and $H_{2}$
$\mathcal{O}\Big((\alpha_{1}\mu_{\max}+\alpha_{2})T+\log(1/\delta)\Big)=\mathcal{O}(\alpha_{H}\mu_{\max}T+\log(1/\delta))$ (2.22)
times (see [ACL2023LCH2, Corollary 16]), where $\alpha_{H}\geq\alpha_{i}$, $i=1,2$.
Applying the block-encoding circuit to the initial input state $|0\rangle_{a^{\prime}}|\tilde{\bm{W}}_{0}\rangle$ gives
$\text{SEL}_{0}|0\rangle_{a^{\prime}}|\tilde{\bm{W}}_{0}\rangle=|0\rangle_{a^{\prime}}\mathcal{U}^{a}(T)|\tilde{\bm{W}}_{0}\rangle+|\bot\rangle,$
where $\mathcal{U}^{a}(T)$ is the approximation of $\mathcal{U}(T)$ and $\tilde{\bm{W}}_{0}=(\Phi^{-1}\otimes I)\bm{W}_{h}(0)$. This step only needs one query to the state preparation oracle $O_{\tilde{w}}$ for $\tilde{\bm{W}}_{0}$.
According to the preceding discussions, we may conclude that there exists a unitary $V_{0}$ such that
$|0^{n_{a}}\rangle|0^{w}\rangle\quad\xrightarrow{V_{0}}\quad\frac{1}{\eta_{0}}|0^{n_{a}}\rangle\otimes\tilde{\bm{W}}_{h}^{a}+|\bot\rangle,$
where $\tilde{\bm{W}}_{h}^{a}$ is the approximate solution of $\tilde{\bm{W}}_{h}$, given by
$\tilde{\bm{W}}_{h}^{a}=\mathcal{U}^{a}(T)\tilde{\bm{W}}_{0}\quad\mbox{and}\quad\eta_{0}=\|\tilde{\bm{W}}_{0}\|=\|\bm{W}_{h}(0)\|\leq C_{e}\|\bm{u}_{I}\|\lesssim\frac{1}{\sqrt{\triangle p}}\sqrt{\|\bm{u}_{0}\|^{2}+T^{2}\|\bm{b}\|_{\mathrm{smax}}^{2}}.$
2.5 An abstract complexity analysis
In this section, we focus on the complexity analysis of the Schrödingerization. According to [JLM24SchrInhom, Theorem 4.4], the error between $\bm{u}_{h}=\mathrm{e}^{p_{k}}(\langle k|\otimes\langle 0|\otimes I)\bm{W}_{h}$ and $\bm{u}$ consists of two parts: one arises from the truncation of the extended domain used for computation in (2.8), and the other results from the spectral discretization in $p$. Suppose $L$ and $R$ are large enough satisfying (2.8), and $\triangle p\simeq\mu_{\max}^{-1}$ is small.
The original Schrödingerization method exhibits first-order convergence in $p$ due to the lack of regularity of $\psi(p)=\mathrm{e}^{-|p|}$ in the initial data in (2.4). Consequently, achieving precision $\epsilon$ may require a small enough mesh size $\triangle p=\mathcal{O}(\epsilon)$. This results in the maximum absolute value among the discrete Fourier modes scaling as $\mathcal{O}(1/\epsilon)$, i.e., $\mu_{\max}=\mathcal{O}(1/\epsilon)$, which is not optimal because the query complexity linearly depends on $\mu_{\max}$ as shown in (2.22).
The parameter $\mu_{\max}$ is proportional to the inverse of the mesh size $\Delta p$, so it should be a function of $\epsilon$. To achieve better precision scaling, the natural idea is to adopt smoother extension of the warped phase transformation $\psi$ so the discrete Fourier transform – which is the spectral method – achieves high order (up to exponential) – accuracy.
Here, we derive the error estimate for the Fourier spectral discretization with smooth initializations, while the detailed proof is given in Section 5. Let $\psi(p)\in H^{r}((-L,R))$ with $\psi^{(k)}(p)\approx 0$ at p=L,Rp=-L,R for $k\leq r$. Then one can apply the discrete Fourier transform to $\psi(p)$. Denote its approximation by $P_{h}\psi$. Noting that $\mu_{\max}=\frac{N_{p}\pi}{L+R}$, the standard approximation estimate [Shenspectral] yields
$\|\psi-P_{h}\psi\|_{L^{2}((-L,R))}\lesssim\Big(\frac{R+L}{N_{p}}\Big)^{r}\|\psi^{(r)}\|_{L^{2}((-L,R))}=\big(\frac{\pi}{\mu_{\max}}\big)^{r}\|\psi^{(r)}\|_{L^{2}((-L,R))}.$ (2.23)
If we assume that the right-hand side of (2.23) is of the same order as $\epsilon$, then
$\mu_{\max}\simeq\pi(1/\epsilon)^{1/r}\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}.$
Theorem 2.1.
Let $\bm{w}(t,p)$ be the exact solution to (2.4), and let $\bm{W}_{h}(t)$ denote the solution of the discrete problem (2.9). Assume that $\psi\in H^{r}(\mathbb{R})$ and decays exponentially on $\mathbb{R}$. Suppose the mesh size $\triangle p$ satisfies
$(\triangle p)^{-1}\simeq\mu_{\max}\simeq\pi(1/\epsilon)^{1/r}\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r},$ (2.24)
where $L$ and $R$ are chosen according to (2.8). Then the following error estimate holds:
$\|\bm{w}(T,p)-\bm{w}_{h}(T,p)\|_{L^{2}((-L,R))}\lesssim\epsilon\|\bm{u}_{I}\|,$ (2.25)
where $\bm{w}_{h}$ is the continuous reconstruction of $\bm{W}_{h}$, given by
$\bm{w}_{h}(t,p)=\sum_{l=0}^{N_{p}-1}\tilde{\bm{w}}_{l,h}(t)\,\phi_{l}(p),\qquad\tilde{\bm{w}}_{l,h}(t)=\frac{1}{N_{p}}\sum_{k=0}^{N_{p}-1}\big((\langle k|\otimes I)\bm{W}_{h}\big)\,\mathrm{e}^{-\mathrm{i}\mu_{l}(p_{k}+L)}.$ (2.26)
For sufficiently large $r$, we have $(1/\epsilon)^{1/r}=\mathcal{O}(1)$. For example, we can assume $\mathrm{e}\leq(1/\epsilon)^{1/r}\leq\mathrm{e}^{2}$ and obtain
$\mu_{\max}\simeq\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}\quad\text{for}\quad\frac{1}{2}\log\frac{1}{\epsilon}\leq r\leq\log\frac{1}{\epsilon}.$ (2.27)
Therefore, if we modify the original function $\psi(p)$ such that
$\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}\lesssim\log(1/\epsilon)\qquad\text{when}\quad r\simeq\log(1/\epsilon),$ (2.28)
then substituting this bound into Eq. (2.22) may imply that the non-unitary dynamic system (2.1) can be simulated from $t=0$ to $t=T$, within an error of $\epsilon$, achieving $\widetilde{\mathcal{O}}(\alpha_{H}T\log(1/\epsilon))$ queries to the HAM-$\text{H}_{\mu}$ oracle.
Eq. (2.28) suggests that we should establish a linear growth of $\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}$ with respect to $r$. However, as will be shown later, a smooth extension of $\psi$ alone is insufficient to achieve such growth and, therefore, cannot yield the optimal cost. This is because we can only derive $\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}\leq Cr^{1/\beta}$ with $\beta\in(0,1)$, which leads to a dependence on $\epsilon$ of $\mathcal{O}(\log^{1/\beta}(1/\epsilon))$. To achieve the optimal convergence rate, one would need $\beta=1$.
The following theorem establishes an abstract framework for the complexity analysis of the Schrödingerization method.
Theorem 2.2.
Let $\varepsilon$ be a positive constant. Suppose that $L$ and $R$ satisfy the condition in (2.8). In addition, we assume that the function $\psi\in H^{r}(\mathbb{R})$ in the initial data of (2.4) decays exponentially on $\mathbb{R}$ and satisfies
$\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}\leq Cr^{1/\beta},\quad\beta\in(0,1],$
where $r\simeq\log(1/\epsilon)$, $C$ is a constant independent of $\epsilon$, and the inverse of the mesh size satisfies (2.24). Here, $\epsilon$ satisfies
$\epsilon\simeq\frac{\epsilon^{\prime}}{(\log(1/\epsilon^{\prime}))^{1/(2\beta)}},\qquad\epsilon^{\prime}=\frac{\varepsilon\|\bm{u}(T)\|}{\|\bm{u}_{I}\|},\qquad\|\bm{u}_{I}\|\simeq\|\bm{u}(0)\|+T\|\bm{b}\|_{\mathrm{smax}},$
with $\|\bm{b}\|_{\mathrm{smax}}$ defined in (2.18). This implies
$\mu_{\max}\lesssim\Big(\log\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}\Big)^{1/\beta}.$
Then, there exists a quantum algorithm that prepares an $\varepsilon$-approximation of the state $|\bm{u}(T)\rangle$ with $\Omega(1)$ success probability and a flag indicating success, using
$\mathcal{O}\Big(\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}\alpha_{H}T\Big(\log\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}\Big)^{1/\beta}\Big)$
queries to the $\text{HAM-T}_{H_{\mu}}$ oracle, where $\alpha_{H}\geq\|H_{i}\|,i=1,2$, and using
$\mathcal{O}\Big(\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}\Big)$
queries to the state preparation oracle for $\tilde{\bm{w}}_{0}$.
Proof.
Let $\bm{W}_{h}(T)$ and $\bm{W}_{h}^{a}(T)$ be the solutions associated with $\mathcal{U}$ and $\mathcal{U}^{a}$, respectively. According to (2.5), one has
$\bm{u}(T)=\mathrm{e}^{p_{k}}(\langle k|\otimes\langle 0|\otimes I)\bm{W}(T),\qquad\bm{u}_{h}^{a}(T)=\mathrm{e}^{p_{k}}(\langle k|\otimes\langle 0|\otimes I)\bm{W}^{a}_{h}(T)$
for some $k\in I_{\Diamond}$, where $\bm{W}(t)=\sum_{ki}w_{i}(t,p_{k})|k,i\rangle$. Here, we can choose $p_{k}=\mathcal{O}(1)$. Then we need to bound the error between $|\bm{u}(T)\rangle$ and $|\bm{u}_{h}^{a}(T)\rangle$. Using the inequality $\|\frac{\bm{x}}{\|\bm{x}\|}-\frac{\bm{y}}{\|\bm{y}\|}\|\leq 2\frac{\|\bm{x}-\bm{y}\|}{\|\bm{x}\|}$ for two vectors $\bm{x},\bm{y}$, we obtain
$\||\bm{u}(T)\rangle-|\bm{u}_{h}^{a}(T)\rangle\|\leq 2\frac{\|\bm{u}(T)-\bm{u}_{h}^{a}(T)\|}{\|\bm{u}(T)\|}.$
This gives
$\|\bm{u}(T)-\bm{u}_{h}^{a}(T)\|\leq\mathrm{e}^{p_{k}}\|\bm{W}(T)-\bm{W}_{h}^{a}(T)\|,\qquad k\in I_{\Diamond}.$
By the triangle inequality,
$\|\bm{W}(T)-\bm{W}_{h}^{a}(T)\|\leq\|\bm{W}(T)-\bm{W}_{h}(T)\|+\|\bm{W}_{h}(T)-\bm{W}_{h}^{a}(T)\|=:\varepsilon_{1}+\varepsilon_{2}.$
For $\varepsilon_{1}$, one has
$\varepsilon_{1}$ $=\|\bm{W}(T)-\bm{W}_{h}(T)\|=(\sum_{k=0}^{N_{p}-1}\|\bm{w}(T,p_{k})-\bm{w}_{h}(T,p_{k})\|^{2})^{1/2}$
$\lesssim\frac{1}{\sqrt{\Delta p}}\|\bm{w}(T,p)-\bm{w}_{h}(T,p)\|_{L^{2}((-L,R))}\lesssim\mu_{\max}^{1/2}\epsilon\|\bm{u}_{I}\|,$
where we have used the estimate (2.25). For $\varepsilon_{2}$, assuming $\|V_{k}^{a}(T)-V_{k}(T)\|\leq\delta$, there holds
$\varepsilon_{2}=\|\bm{W}_{h}(T)-\bm{W}_{h}^{a}(T)\|\leq\|\mathcal{U}-\mathcal{U}^{a}\|\|\bm{W}_{h}(0)\|\leq\delta\|\bm{\psi}\|\|\bm{u}_{I}\|\lesssim\mu_{\max}^{1/2}\delta\|\bm{u}_{I}\|.$
Therefore,
$\||\bm{u}(T)\rangle-|\bm{u}_{h}^{a}(T)\rangle\|\lesssim\mu_{\max}^{1/2}(\epsilon+\delta)\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}.$
The condition on $\psi$ implies that $\mu_{\max}\lesssim\log^{1/\beta}(1/\epsilon)$. Given the above equations, we can require
$\Big(\log\frac{1}{\epsilon}\Big)^{1/(2\beta)}\epsilon\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}\simeq\frac{\varepsilon}{2},\qquad\Big(\log\frac{1}{\epsilon}\Big)^{1/(2\beta)}\delta\frac{\|\bm{u}_{I}\|}{\|\bm{u}(T)\|}\simeq\frac{\varepsilon}{2}.$
From the first equation we can get
$\epsilon\simeq\frac{\varepsilon^{\prime}}{(\log(1/\epsilon^{\prime}))^{1/(2\beta)}},\qquad\epsilon^{\prime}=\frac{\varepsilon\|\bm{u}(T)\|}{\|\bm{u}_{I}\|}.$
This yields
$\mu_{\max}\lesssim\Big(\log\frac{(\log(1/\epsilon^{\prime}))^{1/(2\beta)}}{\varepsilon^{\prime}}\Big)^{1/\beta}\simeq\Big(\log\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}\Big)^{1/\beta},$
$\frac{1}{\delta}\simeq\Big(\log\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}\Big)^{1/\beta}\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}.$
Plugging the above quantities into (2.22), we obtain
$\alpha_{H}T\mu_{\max}+\log\frac{1}{\delta}=\mathcal{O}\Big(\alpha_{H}T\Big(\log\frac{\|\bm{u}_{I}\|}{\varepsilon\|\bm{u}(T)\|}\Big)^{1/\beta}\Big).$
The proof is finished by multiplying the repeated times shown in (2.19). ∎
Remark 2.2.
From this theorem, we observe that the optimal complexity is achieved when $\beta=1$. However, we should relax the assumption that $\psi(p)=\mathrm{e}^{-p}$ for $p>0$, and instead pursue an approximation, as discussed later. Therefore, $\bm{u}_{f}$ is not exactly $\mathrm{e}^{p}\bm{w}$, but rather an approximation of $\bm{u}_{f}$. In Theorem 5.2, we provide an error estimate related to this approximation. Taking this error into account still does not affect our final complexity analysis.
3 Near-optimal dependence with smooth initializations
In this section, we demonstrate that the smooth extension to $p<0$ for $\mathrm{e}^{-p}$ is sufficient to ensure near-optimal dependence on matrix queries. However, as mentioned earlier, the smooth extension alone does not guarantee optimal precision dependence. We present three methods for constructing such a smooth extension.
3.1 The cut-off function
Fig. 2: A smooth extension of $\mathrm{e}^{-p}$
As illustrated in Fig. 2, we seek a smooth function $\psi(p)$ such that $\psi(p)=\mathrm{e}^{-p}$ for $p>0$, and it decays rapidly on the negative axis. This problem can be reduced to finding a smooth approximation $\zeta(p)$ of the step function
$H(p)=\begin{cases}0,&p<0,\\ 1,&p\geq 0,\end{cases}$ (3.1)
and then defining $\psi(p)=\zeta(p)\mathrm{e}^{-p}$. Given that the discrete Fourier transform requires periodic boundary conditions and recognizing the exponential decay of $\mathrm{e}^{-p}$ for positive $p$, we further require $\zeta(p)$ to have compact support. This leads us to consider a cut-off function, which ensures the smoothness of the step function and provides compact support via convolution. The convolution operation is a widely used method for smoothing the step function. It generates various smooth approximations by selecting an appropriate convolution kernel. The cut-off function is a prime example of this approach and has become a fundamental tool in the analysis of partial differential equations. In this subsection, we apply the cut-off technique to construct the desired smooth extension.
We begin by recalling the mollifier, defined as
$\eta(p)=\begin{cases}\frac{1}{C}\exp\left(\frac{1}{|p|^{2}-1}\right),&\quad|p|<1\\ 0,&\quad|p|\geq 1\end{cases},\qquad C=\int_{B_{1}}\exp\left(\frac{1}{|p|^{2}-1}\right)\mathrm{d}p,$ (3.2)
where $B_{1}$ denotes the unit ball in $\mathbb{R}^{n}$ and $C$ is the normalization constant ensuring $\int_{\mathbb{R}^{n}}\eta(p)\,\mathrm{d}p=1$. This function belongs to $C_{0}^{\infty}(\mathbb{R}^{n})$ with support $\overline{B_{1}}$.
Lemma 3.1.
The mollifier satisfies the following estimate for its derivatives in one dimension:
$|\eta^{(k)}(p)|\lesssim C(k):=20^{k}k!\mathrm{e}^{-2k}(2k)^{2k},\quad\forall p\in\mathbb{R}.$ (3.3)
Proof.
We only need to consider $\theta(p)=C\eta(p)$. A direct calculation gives
$\theta^{(k)}(p)=Q_{k}(p)(1-p^{2})^{-2k}\exp\left(\frac{1}{|p|^{2}-1}\right),\qquad|p|<1.$
where $Q_{k}$ is a polynomial and can be recursively defined by
$Q_{0}(p)=1,\quad Q_{k+1}(p)=(1-p^{2})^{2}Q^{\prime}_{k}(p)+2p(2k-1-2kp^{2})Q_{k}(p).$
Let $t=\frac{1}{1-p^{2}}\in(1,\infty)$. For $k\geq 1$, one has
$\theta^{(k)}(p)=Q_{k}(p)\mathrm{e}^{-t}t^{2k},\quad|p|<1,\quad 1<t<\infty.$
Since $\mathrm{e}^{-t}t^{2k}$ achieves the maximum value at $t=2k$ when $t>1$, we have
$|\theta^{(k)}(p)|\leq|Q_{k}(p)|\mathrm{e}^{-2k}(2k)^{2k}.$
It is simple to find that $Q_{k}(p)$ is a polynomial of order $3k$. Let
$Q_{k}(p)=\sum_{j=0}^{3k}a_{k,j}p^{j},\qquad|p|<1,$
and we define the maximum coefficient in absolute value as
$A_{k}=\max_{0\leq j\leq 3k}|a_{k,j}|,\qquad k=1,2,\cdots$
Then it holds
$|Q_{k}(p)|\leq(3k+1)A_{k},\qquad k=1,2,\cdots$
Through careful calculation, it is found that
$(1-p^{2})^{2}Q_{k}^{\prime}(p)$ $=(1-2p^{2}+p^{4})\sum_{j=0}^{3k-1}(j+1)a_{k,j+1}p^{j}=:\sum_{m=0}^{3k+3}S(m)p^{m},$
where $S(m)$ is obtained by collecting powers of $p^{m}$. Similarly,
$2p(2k-1-2kp^{2})Q_{k}(p)=:\sum_{m=0}^{3k+3}T(m)p^{m}.$
By examining the recursive formulas, it is clear that each coefficient satisfies
$|S(m)|\lesssim 12kA_{k},\quad|T(m)|\lesssim 8kA_{k},\quad m\leq 3k+3.$
Hence
$A_{k+1}\leq 20kA_{k},\quad k=1,2,\cdots$
with $A_{1}=2$. By induction, we have
$A_{k}\leq 2\cdot 20^{k-1}(k-1)!.$
Therefore,
$|Q_{k}(p)|\leq(3k+1)A_{k}\leq 20^{k}k!,$
up to a universal multiplicative constant. Combining with the previous estimate yields (3.3).
For any $\varepsilon>0$, we can rescale the function such that its support becomes $\overline{B_{\varepsilon}}$, a closed ball of radius $\varepsilon$. The rescaled function is given by $\eta_{\varepsilon}(p)=\frac{1}{\varepsilon^{n}}\eta\left(\frac{p}{\varepsilon}\right)$. For a function $u\in L_{\text{loc}}^{1}(\Omega)$, the mollifier operator $J_{\varepsilon}$ is defined through convolution as
$J_{\varepsilon}u(p)=(\eta_{\varepsilon}*u)(p)=\int_{\Omega}\eta_{\varepsilon}(p-y)u(y)\mathrm{d}y=\int_{B_{\varepsilon}(p)}\eta_{\varepsilon}(p-y)u(y)\mathrm{d}y,\quad p\in\Omega_{\varepsilon},$ (3.4)
where the domain $\Omega_{\varepsilon}$ is defined by
$\Omega_{\varepsilon}=\left\{p\in\Omega:\overline{B_{\varepsilon}(p)}\subset\Omega\right\}=\left\{p\in\Omega:\text{dist}(p,\partial\Omega)>\varepsilon\right\}.$
It can be verified that $J_{\varepsilon}u\in C^{\infty}(\Omega_{\varepsilon})$ for every $u\in L_{\text{loc}}^{1}(\Omega)$. Furthermore, if $\text{supp}\{u\}\Subset\Omega$, denoting $\delta=\text{dist}(\text{supp}\{u\},\partial\Omega)$, then for $\varepsilon<\delta/4$, we have $J_{\varepsilon}u\in C_{0}^{\infty}(\Omega)$ with $\text{supp}\{J_{\varepsilon}u\}\subset\Omega_{\varepsilon}$.
Fig. 3: A snapshot of the domain of the cut-off function in $\mathbb{R}^{2}$
Now we are ready to describe the construction of the cut-off function with the domain illustrated in Fig. 3.
Lemma 3.2 (cut-off function).
Let $\Omega\subset\mathbb{R}^{n}$ be a non-empty open set, and ${\Omega_{0}}\Subset\Omega$. Define
$\delta=\text{dist}(\Omega_{0},\partial\Omega),\quad d=\frac{\delta}{4},\quad\Omega_{1}=\{p\in\Omega:\text{dist}(p,\Omega_{0})<d\}.$
Let $\phi(p)=\chi_{\Omega_{1}}(p)$ denote the indicator function of $\Omega_{1}$. Then $\zeta=J_{d}{\phi}$ satisfies
$\begin{cases}\zeta\in C_{0}^{\infty}(\Omega),&\text{supp}\{\zeta\}\subset K_{d},\\ \zeta(p)\equiv 1,&p\in\Omega_{0},\\ 0\leq\zeta(p)\leq 1,&p\in\Omega,\end{cases}$
where
${K_{d}}=\{p\in\Omega:\text{dist}(p,\Omega_{1})\leq d\}=\{p\in\Omega:\text{dist}(p,\Omega_{0})\leq 2d\}.$
The function $\zeta$ is referred to as the cut-off function relative to the subset $\Omega_{0}$ in $\Omega$.
The one-dimensional cut-off function satisfies the following estimate for its derivatives [gilbarg2001elliptic, evans2010partial]:
$|\zeta^{(k)}(p)|\lesssim\frac{C(k)}{d^{k}},$ (3.5)
where $C(k)$ is defined in (3.3).
For the Schrödingerization method, we set $\Omega_{0}=(-1,R)$ and $d\geq 1$. Let
$\psi(p)=\zeta(p)\mathrm{e}^{-p}.$ (3.6)
Then it holds that $\text{supp}\{\psi\}\subset[-(1+2d),R+2d]$. The cut-off function and the resulting smooth extension are shown in Fig. 4 for $R=5$ and $d=1$.
Theorem 3.1.
For any $\epsilon>0$, let $d=r\simeq\log(1/\epsilon)$. Then the smooth initialization function $\psi(p)$ defined in (3.6) satisfies
$\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\lesssim r^{2}\simeq\log^{2}(1/\epsilon).$
Proof.
Let $\xi^{(k)}(p)=\zeta^{(k)}(p)\mathrm{e}^{-p}$. Noting that
$\psi^{(r)}(p)=\sum_{k=0}^{r}C_{r}^{k}\zeta^{(k)}(p)(\mathrm{e}^{-p})^{(r-k)}=\sum_{k=0}^{r}(-1)^{r-k}C_{r}^{k}\zeta^{(k)}(p)\mathrm{e}^{-p},$
by careful calculation, for any $p\in\mathbb{R}$, one gets
$|\psi^{(r)}(p)|$ $\leq((C_{r}^{0})^{2}+\cdots+(C_{r}^{r})^{2})^{1/2}(|\xi^{(0)}|^{2}+\cdots+|\xi^{(r)}|^{2})^{1/2}$
$=(C_{2r}^{r})^{1/2}(|\xi^{(0)}|^{2}+\cdots+|\xi^{(r)}|^{2})^{1/2}\leq 2^{r}(|\xi^{(0)}|^{2}+\cdots+|\xi^{(r)}|^{2})^{1/2},$ (3.7)
where the combinatorial equality can be found in [Stanley2011Enumerative]. In addition, we have used the Cauchy-Schwarz inequality and the fact that $C_{2r}^{r}\leq\sum_{k=0}^{2r}C_{2r}^{k}=(1+1)^{2r}=4^{r}$. According to (3.3) and (3.6), if we take $d=r$, then there holds
$|\psi^{(r)}(p)|$ $\leq 2^{r}r^{1/2}\max_{0\leq k\leq r}|\zeta^{(k)}(p)\mathrm{e}^{-p}|\lesssim 2^{r}r^{1/2}\max_{0\leq k\leq r}\frac{C(k)}{d^{k}}\mathrm{e}^{-p}$
$=2^{r}r^{1/2}\max_{0\leq k\leq r}\frac{20^{k}k!\mathrm{e}^{-2k}(2k)^{2k}}{d^{k}}\mathrm{e}^{-p}\leq 40^{r}r^{1/2}(2r)^{2r}.$
The above equation together with $\text{supp}\{\psi\}\subset[-(1+2d),R+2d]$ and $d=r$ yields
$\int_{\mathbb{R}}|\psi^{(r)}(p)|^{2}\mathrm{d}p\lesssim(40^{r}r^{1/2}(2r)^{2r})^{2}\int_{-(1+2d)}^{\infty}\mathrm{e}^{-2p}\mathrm{d}p\lesssim(40^{r}r^{1/2}(2r)^{2r})^{2}\mathrm{e}^{2r}.$
Therefore, by taking $r\simeq\log(1/\epsilon)$, we obtain
$\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\lesssim(40^{r}r^{1/2}(2r)^{2r})^{1/r}\lesssim r^{2}\lesssim\log^{2}(1/\epsilon).$
This completes the proof. ∎
According to Theorem 2.2, the above result implies a sub-optimal precision dependence. Ref. [Johnson15bump] states that the mollifier $\eta(p)$ decays in the Fourier domain asymptotically as $\hat{\eta}(w)=(|w|^{-3/4})\mathrm{e}^{-\sqrt{|w|}}$. This exhibits super-polynomial decay, as the exponent involves a square root of $|w|$, rather than exponential decay. This corresponds to the case $\beta=1/2$ in [ACL2023LCH2], which consequently leads to $\mathcal{O}(\log^{2}(1/\epsilon))$ for the LCHS method in terms of precision.
Fig. 4: The cut-off function and the resulting smooth extension
3.2 Higher-order interpolation
An alternative approach to constructing the smooth function in the extended domain is to utilize high-order interpolations. To do so, we rewrite the initial function as
$\psi(p)=\begin{cases}h(p),&p\in(-\infty,0],\\ \mathrm{e}^{-p},&p\in(0,\infty),\end{cases}$
where $h(p)$ is defined by
$h(p)=\mathcal{P}_{2r-1}(p),\quad p\in[-1,0],\quad\quad h(p)=\mathrm{e}^{p},\quad p\in(-\infty,-1).$ (3.8)
Here, $\mathcal{P}_{2r-1}(p)$ is a Hermite interpolation polynomial whose degree does not exceed (2r1)(2r-1) [SBBGW80, Section 2.1.5], satisfying
$\big(\partial_{p^{k}}\mathcal{P}_{2r-1}(p)\big)|_{p=0}$ $=\big(\partial_{p^{k}}(\mathrm{e}^{-p})\big)|_{p=0}=(-1)^{k},$
$\big(\partial_{p^{k}}\mathcal{P}_{2r-1}(p)\big)|_{p=-1}$ $=\big(\partial_{p^{k}}(\mathrm{e}^{p})\big)|_{p=-1}=\mathrm{e}^{-1},$
where $0\leq k\leq r-1$ is an integer. It is simple to check that $\psi\in C^{r-1}(\mathbb{R})$ and $\psi(p)\in H^{r}((-L,R))$ after restricting the extended domain to a limited interval. The explicit formula of $\mathcal{P}_{2r-1}(p)$ is given by
$\mathcal{P}_{2r-1}(p)=\sum_{k=0}^{r-1}\big(\mathrm{e}^{-1}L_{0k}(p)+(-1)^{k}L_{1k}(p)\big),$
where $L_{ik}$ are generalized Lagrange polynomials defined recursively for k=r2k=r-2, r3r-3, $\cdots$, 0,
$L_{0k}(p):=l_{0k}(p)-\sum_{\nu=k+1}^{r-1}l_{0k}^{(\nu)}(-1)L_{0\nu}(p),\quad L_{1k}(p):=l_{1k}(p)-\sum_{\nu=k+1}^{r-1}l_{1k}^{(\nu)}(0)L_{1\nu}(p),$
with the starting polynomial for k=r1k=r-1
$L_{0r-1}(p):=l_{0r-1}(p),\qquad L_{1r-1}(p):=l_{1r-1}(p).$
The auxiliary polynomials are
$l_{0k}(p):=\frac{(-1)^{r}(p+1)^{k}p^{r}}{k!},\qquad l_{1k}(p):=\frac{p^{k}(p+1)^{r}}{k!}.$
According to the discussion in [JLM24SchrBackward], the target variable $\bm{u}(t)=\mathrm{e}^{p}\bm{v}(t,p)$ still holds for all $p>0$, since we do not care the solution when $p<0$. We provide the snapshots of $\psi$ for $r=2,4,6,8,10$ in Fig. 5.
Fig. 5: The smooth initial data of $\psi(p)$ by using high-order interpolation.
By employing the mollifier technique described in Section 3.1, we can identify a smooth function $\varphi\in C^{\infty}(\mathbb{R})$ such that $\varphi(p)\big|_{(-\infty,-1)\cup(0,\infty)}=\mathrm{e}^{-|p|}$. Consequently, $\mathcal{P}_{2r-1}(p)$ can be interpreted as an interpolation of $\varphi$. Given that $\|\varphi^{(r)}\|_{L^{\infty}(-1,0)}$ is bounded, the $H^{r}$ norm of $\psi$ over the finite interval [L,R][-L,R] remains bounded. Unfortunately, while $\psi$ is contained in $H^{r}$ with any fixed $r\geq 1$, its exclusion from the $C^{\infty}$-class necessarily restricts the achievable complexity to a quasi-optimal order. For the Schrödingerization method, the truncation of the extended domain must satisfy (2.8), it follows that $\bm{w}(t,-L)\approx\bm{w}(t,R)\approx 0$ [JLM24SchrInhom]. Therefore, spectral methods can be used.
3.3 Fourier transform
Building upon similar principles and utilizing the continuous Fourier transform in $p$, [ALL2023LCH] introduces an algorithm for implementing Linear Combination of Hamiltonian Simulation (LCHS). The original LCHS approach in [ALL2023LCH] is also a first-order method due to the slow decay rate of the integrand as a function of $k$, where $k$ is the continuous Fourier mode. In the continuous scenario, the integrand function with respect to $k$ is the Fourier transform of $\psi(p)=\mathrm{e}^{-|p|}$, given by $\frac{1}{\pi(1+k^{2})}$. It decays only quadratically, necessitating the truncation interval choice of [K,K][-K,K] with $K=\log(1/\varepsilon)$. This introduces a computational overhead of $\mathcal{O}(1/\varepsilon)$, as $kH_{1}(s)-H_{2}(s)$ may have a spectral norm as large as $K\|H_{1}(s)\|$. This limitation was addressed in [ACL2023LCH2] by replacing the original integrand with a new kernel function decaying at a near-exponential rate $\mathrm{e}^{-c|k|^{\beta}}$, where $\beta\in(0,1)$. Consequently, they no longer need to truncate the interval at $K=\mathcal{O}(1/\varepsilon)$, and instead use the much smaller cutoff $K=(\log(1/\varepsilon))^{1/\beta}$. The improved LCHS method requires
$\widetilde{\mathcal{O}}\left(\frac{\|\bm{u}(0)\|+\|\bm{b}\|_{L^{1}}}{\|\bm{u}(T)\|}\alpha_{H}T(\log(1/\varepsilon))^{\gamma}\right)$
queries to the HAM-T oracle, where $\|\bm{b}\|_{L^{1}}=\int_{0}^{T}\|\bm{b}(s)\|\mathrm{d}s$, and $\gamma=1+1/\beta$ and $\beta$ for linear systems with time-dependent and time-independent coefficients, respectively. This leads to an exponential reduction in the Hamiltonian simulation time with respect to $\varepsilon$ compared to the original LCHS. Since $\beta\in(0,1)$, it holds that $1/\beta>1$, indicating sub-optimal behavior with respect to queries to the HAM-T oracle.
Fig. 6: Fourier transforms of the kernel functions (see Figure 3 of [ACL2023LCH2])
We emphasize that the fundamental principle remains consistent with our approach with smooth initialization. Indeed, as illustrated in Fig. 6, the Fourier transforms of their kernel functions are $\mathrm{e}^{-x}$ for $x>0$, whereas they exhibit significant differences on the negative real axis. Therefore, the transformed kernel function can be interpreted as a smooth extension in the $p$ space.
4 Establishing the optimal precision dependence
In this section, we demonstrate how to achieve optimal dependence on matrix queries by choosing suitable smooth initializations.
4.1 Conditions for optimal precision dependence
According to the discussion in Section 2.5, if we require the original function $\psi(p)$ to satisfy
$\|\psi^{(r)}\|_{L^{2}((-L,R))}^{1/r}\lesssim\log(1/\epsilon)\qquad\text{when}\quad r\simeq\log(1/\epsilon),$ (4.1)
then substituting this bound into Theorem 2.2 implies that the non-unitary dynamic system (2.1) can be simulated from $t=0$ to $t=T$, within an error of $\varepsilon$, achieving $\widetilde{\mathcal{O}}(\alpha_{H}T\log(1/\epsilon))$ queries to the HAM-$\text{H}_{\mu}$ oracle. This achieves optimal dependence on matrix queries.
However, it is impossible to achieve the bound in (4.1) while still requiring that $\psi(p)=\mathrm{e}^{-p}$ for $p\geq 0$. The reason is as follows. Let $\hat{\psi}(w)$ denote the inverse Fourier transform of $\psi(p)$, defined by
$\hat{\psi}(w)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\psi(p)\mathrm{e}^{\mathrm{i}wp}\mathrm{d}p.$
It follows from the Parseval’s identity that
$\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}=\|w^{r}\hat{\psi}\|_{L^{2}(\mathbb{R})}.$ (4.2)
  • If $\psi(p)$ is $C^{\infty}$ but not analytic, one can only expect super-polynomial decay of $\hat{\psi}$:
    $|\hat{\psi}(w)|\leq C\mathrm{e}^{-c|w|^{\beta}},\qquad 0<\beta<1.$
    For example, the mollifier $\eta(p)$ in (3.2) decays in the Fourier domain asymptotically as $\hat{\eta}(w)=(|w|^{-3/4})\mathrm{e}^{-\sqrt{|w|}}$ [Johnson15bump]. One can show that in this case,
    $\|w^{r}\hat{\psi}\|_{L^{2}(\mathbb{R})}\leq Cr^{r/\beta},$
    which implies
    $\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\leq Cr^{1/\beta}=\mathcal{O}(\log^{1/\beta}(1/\epsilon))$
    when $r\simeq\log(1/\epsilon)$.
  • If $\psi(p)$ is analytic, according to the Paley-Wiener theorem, we have the exponential decay of $\hat{\psi}$:
    $|\hat{\psi}(w)|\leq C\mathrm{e}^{-c|w|},$
    which implies the desired bound
    $\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}=\mathcal{O}(r)=\mathcal{O}(\log(1/\epsilon))$
    when $r\simeq\log(1/\epsilon)$.
As demonstrated in [ACL2023LCH2], the LCHS method cannot achieve optimal scaling with respect to the error tolerance $\varepsilon$. The underlying reason is that $\hat{\psi}(w)$ cannot exhibit exponential decay under the requirement that $\psi(p)=\mathrm{e}^{-p}$ for $p\geq 0$. The Schrödingerization and LCHS frameworks are closely related, as they share similar foundational mathematical principles. This connection further implies that the Schrödingerization method also attains only suboptimal scaling in $\epsilon$.
Based on the above discussion, to achieve optimal dependence on precision, we should relax the condition that $\psi(p)=\mathrm{e}^{-p}$ for $p\geq 0$ and instead seek an approximate version. The required assumptions on $\psi(p)$ are as follows:
Conditions on $\psi(p)$
  • (H1)
    $\psi(p)$ exhibits exponential decay on $\mathbb{R}$ such that
    $\psi(p)\lesssim\mathrm{e}^{-|p|}\leq 2\epsilon,\qquad\;p\in(-\infty,-L+\lambda_{\max}^{-}(H_{1})T)\cup(R-\lambda_{\max}^{+}(H_{1})T,+\infty),$
    where $L$ and $R$ satisfy (2.8).
  • (H2)
    For $p\in[p_{*},R]$, the condition $|\psi(p)-\mathrm{e}^{-p}|\leq\epsilon$ holds, where $p_{*}\leq 1/2$,
  • (H3)
    There exist a constant $C$, independent of $r$ and $\epsilon$, such that
    $\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\leq Cr,\qquad r\simeq\log(1/\epsilon).$ (4.3)
The first assumption (H1) on $\psi$ is required to approximate the infinite domain problem by a periodic problem on the truncated interval [L,R][-L,R], with specific error estimates provided in Theorem 5.3. The second assumption ensures the recovery of the target variable from the warped phase transformation in the Schrödinger-type formulation. It is noteworthy that if $\psi(p)=\mathrm{e}^{-p}$, the target variable can be recovered without error (see Theorem 5.1). If $\psi(p)$ satisfies an approximate condition (H2) , the corresponding error estimates are given in Theorem 5.2. The third assumption (H3) is introduced to derive the optimal estimate for $\mu_{\max}$ , with the main result presented in Theorem 2.2.
4.2 Construction of the function
Motivated by the construction in the cut-off function method (cf. (3.6)), we define
$\psi(p)=\phi(p)\mathrm{e}^{-p},$
where the function $\phi(p)$ decays super-exponentially as $p\to$-$\infty$; that is, $\lim\limits_{p\to-\infty}\phi(p)\mathrm{e}^{-p}=0$, and satisfies $|\phi(p)-1|\leq\epsilon$ for $p\geq p_{*}$. Moreover, $\phi$ fulfills the following norm constraint
$\|\phi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\leq Cr,\qquad r\geq 1.$ (4.4)
As analyzed in Section 4.1, the suboptimal cost of the cut-off method discussed in Section 3.1 is primarily due to the lack of analyticity of the convolution kernel in (3.4). Specifically, the mollifier is not analytic, meaning it cannot be expressed as a convergent power series. The motivation for using the mollifier as the convolution kernel is that it provides a smooth function with compact support, thus realizing the exact periodic boundary conditions. However, if we relax this condition and seek a function that approximately satisfies the periodic boundary conditions, the most natural choice is the Gaussian $\mathrm{e}^{-p^{2}}$. In this case, a careful calculation reveals that
$\phi(p):=(H(t)*\mathrm{e}^{-t^{2}})(p)=\int_{\mathbb{R}}H(t)\mathrm{e}^{-(p-t)^{2}}\mathrm{d}t=\frac{\text{erf}(p)+1}{2},$ (4.5)
where $H(t)$ is the step function in (3.1), and $\text{erf}(p)$ is the error function, which is defined as
$\text{erf}(p)=\frac{2}{\sqrt{\pi}}\int_{0}^{p}\mathrm{e}^{-t^{2}}\mathrm{d}t.$
We note that the error function is also used in [Low25LCHS] to construct the optimal LCHS method, and our work is partially inspired by the approach presented in [Low25LCHS]. For later uses, we also introduce the complementary error function $\text{erfc}(p)=1-\text{erf}(p)$.
Fig. 7: Error function $\text{erf}(p)$
As shown in Fig. 7, the error function is an odd function that approaches $\pm 1$ at a super-exponential rate (see Eq. (4.7)). This implies that $\phi(p)$ in (4.6) tends to 1 at a super-exponential rate. However, this does not imply that $\psi(p)\approx 1$ for $p\geq p_{*}$, where $0\leq p_{*}\leq\frac{1}{2}$. To address this, we rescale the error function and define
$\phi(p)=\frac{\text{erf}(ap)+1}{2},$ (4.6)
where $a\geq 1$ is a constant to be determined. Setting $x=ap$, we aim to find the lower bound of $x$ such that
$\Big|\frac{\text{erf}(x)+1}{2}-1\Big|=\Big|-\frac{1}{2}\text{erfc}(x)\Big|\leq\epsilon.$
For fixed $x>0$, we have the following expansion
$\text{erfc}(x)=\frac{\mathrm{e}^{-x^{2}}}{x\sqrt{\pi}}\Big(1-\frac{1}{2x^{2}}+\frac{3}{4x^{2}}-\frac{15}{8x^{6}}+\cdots\Big),$
which is an alternating series, leading to
$\text{erfc}(x)\leq\frac{\mathrm{e}^{-x^{2}}}{x\sqrt{\pi}}\leq\mathrm{e}^{-x^{2}},\qquad x\geq 1.$ (4.7)
This implies
$ap=x\geq\log^{1/2}\frac{1}{\epsilon},$
and thus, if we choose
$a=2\log^{1/2}\frac{1}{\epsilon},$
we obtain $p_{*}=1/2$. Unless otherwise specified, we set $p_{*}=1/2$ throughout the following analysis.
Remark 4.1.
Since the error function is odd and asymptotically approaches $\pm 1$, it follows that $\phi(p)$ decays to zero on the negative axis at the same rate as shown in (4.7). This implies that $\psi(p)$ decays super-exponentially as $p\to$-$\infty$ and exponentially as $p\to+\infty$.
Theorem 4.1.
Let $\phi(p)$ be defined in (4.6) with $a=2\log^{1/2}(1/\epsilon)$. Then the function $\psi(p)=\phi(p)\mathrm{e}^{-p}$ satisfies the required conditions (H1)–(H3).
Proof.
We only need to verify the inequality (4.3). Let $\xi^{(k)}(p)=\phi^{(k)}(p)\mathrm{e}^{-p}$. Following the same calculation in (3.7), one gets
$|\psi^{(r)}(p)|\leq 2^{r}(|\xi^{(0)}|^{2}+\cdots+|\xi^{(r)}|^{2})^{1/2}.$
This implies
$|\psi^{(r)}(p)|^{1/r}\lesssim\max_{0\leq k\leq r}|\xi^{(k)}|^{1/r}=\max_{0\leq k\leq r}|\phi^{(k)}(p)\mathrm{e}^{-p}|^{1/r}\lesssim a\max_{0\leq k\leq r}|\text{erf}^{(k)}(ap)\mathrm{e}^{-p}|^{1/r},$
since $\phi^{(k)}(p)=\frac{1}{2}a^{k}\text{erf}^{(k)}(ap)$.
A direct calculation gives
$\text{erf}^{(k)}(ap)=\frac{2}{\sqrt{\pi}}(-1)^{k-1}H_{k-1}(ap)\mathrm{e}^{-a^{2}p^{2}},\qquad 1\leq k\leq r,$
where $H_{k}$ is the Hermitian polynomial, defined by $H_{k}(p)=(-1)^{k}\mathrm{e}^{p^{2}}(\mathrm{e}^{-p^{2}})^{(k)}$, which leads to
$|\text{erf}^{(k)}(ap)\mathrm{e}^{-p}|\lesssim|H_{k-1}(ap)\mathrm{e}^{-a^{2}p^{2}-p}|.$
It is known that
$|H_{k}(ap)|\leq C2^{k/2}\sqrt{k!}\mathrm{e}^{a^{2}p^{2}/2},$
where $C\approx 1.086435$ (see [Shenspectral, Eq. (7.66)] for example). Therefore, we obtain
$\|\text{erf}^{(k)}(ap)\mathrm{e}^{-p}\|_{L^{2}(\mathbb{R})}^{1/r}\lesssim 2^{\frac{k}{2r}}(k!)^{1/(2r)}\|\mathrm{e}^{-a^{2}p^{2}/2-p}\|_{L^{2}(\mathbb{R})}\leq r^{1/2},\quad k\leq r,$
which gives
$\|\psi^{(r)}\|_{L^{2}(\mathbb{R})}^{1/r}\lesssim ar^{1/2}.$
The proof is completed by noting that $a\simeq r^{1/2}\simeq\log^{1/2}(1/\epsilon)$. ∎
Remark 4.2.
Among the three conditions, the most challenging to satisfy is (H3). In fact, functions satisfying (H3) belong to the Gevrey class of order 1. Recall that a function $f\in C^{\infty}(\mathbb{R})$ is said to be in the Gevrey class $G^{s}(\mathbb{R})$ for $s\geq 1$ if, for every compact set $K\subset\mathbb{R}$, there exist constants $C,R>0$ (independent of $r$) such that
$\sup_{p\in\mathbb{R}}|f^{(r)}(p)|\leq CR^{r}(r!)^{s}\qquad\forall r\in\mathbb{N}.$
In particular, the case $s=1$ corresponds to the Gevrey-1 class, which corresponds to the condition (H3).
This definition implies that the construction of $\psi$ is not unique. For example, define
$\chi(p)=\big(\int_{-\infty}^{\infty}\mathrm{e}^{-t^{4}}\mathrm{d}t\big)^{-1}\int_{0}^{p}\mathrm{e}^{-t^{4}}\mathrm{d}t,\quad\text{and}\quad\tilde{\phi}(p)=\chi(ap)+\frac{1}{2},$
where $a$ is an adjustable parameter chosen so that (H2) is satisfied. Following the same line of the proof, we obtain a new initial function $\psi(p)=\tilde{\phi}(p)\mathrm{e}^{-p}$.
Remark 4.3.
When there is no inhomogeneous term, the introduction of auxiliary constant variables is unnecessary. In this case, the total query complexity for $H_{1}$ and $H_{2}$ is $\tilde{\mathcal{O}}\big(\frac{\|\bm{u}(0)\|}{\|\bm{u}(T)\|}\alpha_{H}T\mu_{\max}\big)$. When the system is time-dependent, we employ the approach described in Section 2.2 to obtain a time-independent Hamiltonian system by introducing an additional dimension. Following the analysis in Theorem 2.2, it is straightforward to derive that the query complexity for the time-dependent system is $\tilde{\mathcal{O}}\big(\frac{\|\bm{u}(0)\|}{\|\bm{u}(T)\|}\alpha_{H}T\mu_{\max}\mu_{\max}^{s}\big)$, where $\mu_{\max}^{s}=\max_{l\in[N_{s}]}|\mu_{l}^{s}|$. After applying the same smoothing technique, the query complexity approaches $\widetilde{\mathcal{O}}(\frac{\|\bm{u}(0)\|}{\|\bm{u}(T)\|}\alpha_{H}T(\log\frac{1}{\varepsilon})^{2})$, which includes an additional logarithmic term with respect to $\varepsilon$, consistent with the observation in [ACL2023LCH2] (see Table 1 there).
Remark 4.4.
Recent results in [ALWZ25] establish quantum query lower bounds for solving ODEs by reduction to quantum linear system solvers. These bounds show that any algorithm requires at least $\Omega(T^{\alpha})$ queries for sufficiently long evolution time $T$ and at least $\Omega(\log^{\alpha}(1/\varepsilon))$ queries for target precision $\varepsilon$ in the worst case, where $\alpha$ is the query lower bound exponent for quantum linear system solvers. According to [Costa2021QLSA], one has $\alpha=1$. This implies that our Schrödingerization algorithm achieves optimal dependence on matrix queries.
5 Error estimate for the Schrödingerization method
In this section, we analyze the error of the Schrödingerization method with smooth initial functions. The discretization error consists of two parts: recovering the target variable $\bm{u}_{f}$, the solution of (2.2), from $\bm{w}$, the solution of (2.4); and the numerical error from applying the Fourier spectral method to (2.4). For simplicity, we assume that both $A$ and the source term $\bm{b}$ are time-independent; the time-dependent case can be treated analogously.
5.1 Recovery from (2.4) with smooth initializations
In the following, we provide a rigorous basis for the recovery of $\bm{u}_{f}(t)$ from $\bm{w}(t,p)$ under the stated conditions.
Lemma 5.1.
Let $\bm{\omega}:[0,T]\times\mathbb{R}\to\mathbb{C}^{N}$ be a solution to
$\frac{\mathrm{d}\bm{\omega}}{\mathrm{d}t}=-H_{1}\partial_{p}\bm{\omega}+\mathrm{i}H_{2}\bm{\omega},\quad\bm{\omega}(0,p)=\begin{cases}\textbf{0},\quad p\in[p_{*},R]\\ \bm{\omega}_{0}(p),\quad p\in\mathbb{R}\backslash[p_{*},R],\end{cases}$
where $\bm{\omega}_{0}(p)$ is a given function. Then we have
$\bm{\omega}(T,p)=\bm{0},\qquad p\in(p_{*}+\lambda_{\max}^{+}(H_{1})T,~~R-\lambda_{\max}^{-}(H_{1})T).$
Proof.
We first derive a local conservation law for the squared amplitude. Since $H_{1},H_{2}$ are Hermitian, from the evolution equation $\partial_{t}\bm{\omega}=-H_{1}\partial_{p}\bm{\omega}+\mathrm{i}H_{2}\bm{\omega}$, we obtain
$\partial_{t}\big(\bm{\omega}^{\dagger}\bm{\omega}\big)=\bm{\omega}^{\dagger}\partial_{t}\bm{\omega}+(\partial_{t}\bm{\omega})^{\dagger}\bm{\omega}=-\partial_{p}\big(\bm{\omega}^{\dagger}H_{1}\bm{\omega}\big).$ (5.1)
For $t\in[0,T]$, define the moving interval
$I(t)=\big(p_{*}+\lambda_{\max}^{+}(H_{1})t,\;\;R-\lambda_{\max}^{-}(H_{1})t\big),$
and the localized energy
$E_{\omega}(t):=\int_{I(t)}\|\bm{\omega}(t,p)\|^{2}\,\mathrm{d}p.$
By the Leibniz’s rule and (5.1), one finds
$\frac{\mathrm{d}}{\mathrm{d}t}E_{\omega}(t)=$ $-\Big[\,\bm{\omega}^{\dagger}H_{1}\bm{\omega}\,\Big]_{p=p_{*}+\lambda_{\max}^{+}(H_{1})t}^{R-\lambda_{\max}^{-}(H_{1})t}+\lambda_{\max}^{-}(H_{1})\|\bm{\omega}(t,R-\lambda_{\max}^{-}(H_{1})t)\|^{2}$
$\quad-\lambda_{\max}^{+}(H_{1})\|\bm{\omega}(t,p_{*}+\lambda_{\max}^{+}(H_{1})t)\|^{2}\leq 0.$
At $t=0$, the initial condition ensures $\bm{\omega}(0,p)\equiv\bm{0}$ for $p\in(p_{*},R)$, hence $E(0)=0$. By monotonicity, we conclude $E_{\omega}(t)\equiv 0$ for all $t\in[0,T]$. Therefore, $\bm{\omega}(t,p)\equiv 0$ almost everywhere in $I(t)$, and by continuity, it follows that $\bm{\omega}(T,p)\equiv\bm{0}$ for $p\in\big(p_{*}+\lambda_{\max}^{+}(H_{1})T,\;\;R-\lambda_{\max}^{-}(H_{1})T\big)$. ∎
Remark 5.1.
Following the proof, it is clear that if $\bm{\omega}(0,p)=\epsilon\bm{u}_{I}$ for $p\in[p_{*},R]$, then
$\int_{p_{*}+\lambda_{\max}^{+}(H_{1})T}^{R-\lambda_{\max}^{-}(H_{1})T}\|\bm{\omega}(T,p)\|^{2}\mathrm{d}p\leq\int_{p_{*}}^{R}\|\bm{\omega}(0,p)\|^{2}\mathrm{d}p=\epsilon^{2}\|\bm{u}_{I}\|^{2}(R-p_{*}).$
Theorem 5.1.
Assume $\psi(p)\in L^{2}(\mathbb{R})$ with $\psi(p)=\mathrm{e}^{-p}$ in $(0,R)$, and $R>(\lambda_{\max}^{-}(H_{1})+\lambda_{\max}^{+}(H_{1}))T$. Then the solution of (2.1) can be recovered by
$\bm{u}_{f}(T)=\mathrm{e}^{p}\bm{w}(T,p),\quad p\in(\lambda_{\max}^{+}(H_{1})T,~~R-\lambda_{\max}^{-}(H_{1})T),$
where $\bm{w}(T,p)$ is the solution to (2.4).
Proof.
The function $\widetilde{\bm{w}}(t,p)=\mathrm{e}^{-p}\mathrm{e}^{(H_{1}+\mathrm{i}H_{2})t}\bm{u}_{0}$ satisfies both the PDE and the initial condition whenever the backward characteristics from $(t,p)$ remain inside $(0,R)$, i.e., $p-\lambda_{j}t\in(0,R)$, where $\lambda_{j}$ is the eigenvalue of $H_{1}$. This requires $\lambda_{\max}^{+}(H_{1})T<p<R-\lambda_{\max}^{-}(H_{1})T$. Define the error vector $\bm{e}_{w}=\bm{w}-\widetilde{\bm{w}}$. Then $\bm{e}_{w}(t,p)$ satisfies the same PDE, and $\bm{e}_{w}(0,p)\equiv\bm{0}$ for all $p\in(0,R)$. The proof is completed by applying Lemma 5.1. ∎
Theorem 5.2.
Assume that $\psi(p)\in L^{2}(\mathbb{R})$ satisfies condition (H2). The recovery from $\bm{w}(T,p)$, which is the solution to (2.4), is defined by
$\bm{u}_{f}^{*}(T,p)=\mathrm{e}^{p}\bm{w}(T,p)\qquad p\in(p_{*}+\lambda_{\max}^{+}(H_{1})T,~~R-\lambda_{\max}^{-}(H_{1})T).$
Assume $p_{*}=\frac{1}{2}$ and $R=\lambda_{\max}^{-}(H_{1})T+\mathcal{O}(1)$. The $L^{2}$ error estimate between the recovery and the solution to (2.1) is given by
$\frac{1}{L_{*}}\int_{p_{*}+\lambda_{\max}^{+}(H_{1})T}^{R-\lambda_{\max}^{-}(H_{1})T}\|\bm{u}_{f}^{*}(T,p)-\bm{u}_{f}(T)\|^{2}\mathrm{d}p\lesssim\epsilon^{2}\|\bm{u}_{I}\|^{2},$
where $L_{*}=R-p_{*}-(\lambda_{\max}^{+}(H_{1})T+\lambda_{\max}^{-}(H_{1})T)$ is the length of the recovery domain.
Proof.
Since the matrix $A$ and the vector $\bm{b}$ in system (2.1) are time-independent and satisfy $\lambda(A+A^{\dagger})\leq 0$, it follows that $\lambda_{\max}^{+}(H_{1})T\leq 1/2$. The proof is completed by applying Theorem 5.1 and Remark 5.1. ∎
5.2 Error estimate of spectral discretization
In this subsection, we provide the detailed proof of Theorem 2.1. The error associated with (2.25) arises from two sources: (i) the truncation error introduced by restricting the equation for $\bm{w}$ from the whole space to a finite domain with periodic boundary conditions, and (ii) the discretization error of the Fourier spectral method applied to the periodic problem on the finite domain.
5.2.1 Error estimate of the truncation
The vector $\bm{w}(\cdot,p)$ for Eq. (2.4) is defined in $\mathbb{R}$. However, in the implementation, $\mathbb{R}$ is truncated to the interval [L,R][-L,R], with $L,R$ satisfying (2.8), resulting in the solution $\bm{w}_{h}(\cdot,p)$ in (2.9). In the following, we will characterize the truncation error associated with this approximation.
It is apparent that the discretization (2.9) serves as an approximation for the following system with periodic boundary conditions:
$\begin{cases}\frac{\partial}{\partial t}\mathcal{W}=-H_{1}\partial_{p}\mathcal{W}+\mathrm{i}H_{2}\mathcal{W},\quad 0<t<T,~~-L<p<R,\\ \mathcal{W}(t,-L)=\mathcal{W}(t,R),\\ \mathcal{W}(0,p)=\psi(p)\bm{u}_{I},\end{cases}$ (5.2)
where $\psi(p)$ satisfies $\textbf{(H1)}-\textbf{(H2)}$, and $\psi\in H^{r}((-L,R))$. Noting that $\psi^{(k)}(p)\approx 0$ at p=L,Rp=-L,R for $k\leq r$, it implies that each entry of $\mathcal{W}(0,p)$ can be treated as a function in $H_{p}^{k}[-L,R]$ for any $0\leq k\leq r$, which consists of functions with derivatives of order up to (k1)(k-1) being periodic on [L,R][-L,R]. It is important to note that this estimate applies not only to the initial data but also to the solution of (5.2), since $\mathcal{W}$ satisfies a transport equation in the $p$ direction thus preserves the regularity in $p$ in the initial data as time evolves.
We observe that to estimate the error between $\bm{w}_{h}$ and $\bm{w}$, it suffices to bound the error between $\mathcal{W}$ and $\bm{w}$.
Lemma 5.2.
Let $\bm{w}:[0,T]\times\mathbb{R}\to\mathbb{C}^{N}$ be the solution to (2.4) and $\psi(p)$ satisfies (H1). Then it holds
$\left|\int_{0}^{T}\bm{w}^{\dagger}(t,-L)H_{1}\bm{w}(t,-L)\mathrm{d}t\right|$ $\lesssim\,\int_{-\infty}^{-L+\lambda_{\max}^{-}(H_{1})T}\|\bm{w}(0,p)\|^{2}\mathrm{d}p,$ (5.3)
$\left|\int_{0}^{T}\bm{w}^{\dagger}(t,R)H_{1}\bm{w}(t,R)\mathrm{d}t\right|$ $\lesssim\,\int_{R-\lambda_{\max}^{+}(H_{1})T}^{+\infty}\|\bm{w}(0,p)\|^{2}\mathrm{d}p.$ (5.4)
Proof.
Define the local energy density $E(t,p)=\|\bm{w}(t,p)\|^{2}=\bm{w}(t,p)^{\dagger}\bm{w}(t,p)$. Multiplying (2.4) by $\bm{w}^{\dagger}$ from the left and using the Hermiticity of $H_{2}$, we obtain the conservation law
$\partial_{t}E(t,p)+\partial_{p}\big(\bm{w}^{\dagger}H_{1}\bm{w}\big)=0.$ (5.5)
Let $\{\chi_{n}\}_{n\geq 1}\subset C_{c}^{\infty}(\mathbb{R})$ be a sequence of smooth cutoff functions such that $\chi_{n}(x)=1$ for $x\leq 0$, $\chi_{n}(x)=0$ for $x\geq 1/n$, and $\chi_{n}^{\prime}(x)\leq 0$. For fixed $t\in[0,T]$, define the moving cutoff
$\Phi_{n,s}(p):=\chi_{n}\big(p+L-\lambda_{\max}^{-}(H_{1})(t-s)\big),\qquad s\in[0,t],$
and the localized energy
$I_{n}(s):=\int_{\mathbb{R}}\Phi_{n,s}(p)\,E(s,p)\,\mathrm{d}p.$
Differentiating $I_{n}(s)$ with respect to $s$, and using (5.5) together with integration by parts, one obtains
$I_{n}^{\prime}(s)=\int_{\mathbb{R}}\chi_{n}^{\prime}\left(p+L-\lambda_{\max}^{-}(H_{1})(t-s)\right)\,\big(\lambda_{\max}^{-}(H_{1})E(s,p)+\bm{w}^{\dagger}H_{1}\bm{w}(s,p)\big)\,\mathrm{d}p.$
The boundary terms from integration by parts vanish–after ignoring the error of $O(\epsilon)$– due to the exponential decay of $\bm{w}$. Since $\chi_{n}^{\prime}\leq 0$ and $\lambda_{\max}^{-}(H_{1})E\leq\bm{w}^{\dagger}H_{1}\bm{w}\leq\lambda_{\max}^{+}(H_{1})E$, the integrand is positive, hence $I_{n}^{\prime}(s)\leq 0$ for all $s\in[0,t]$. Therefore $I_{n}(t)\leq I_{n}(0)$.
By the construction,
$I_{n}(t)=\int_{\mathbb{R}}\chi_{n}(p+L)\,\|\bm{w}(t,p)\|^{2}\mathrm{d}p,\qquad I_{n}(0)=\int_{\mathbb{R}}\chi_{n}(p+L-\lambda_{\max}^{-}(H_{1})t))\,\|\bm{w}(0,p)\|^{2}\mathrm{d}p.$
Taking $n\to\infty$ , we obtain
$\int_{-\infty}^{-L}\|\bm{w}(t,p)\|^{2}\mathrm{d}p\;\leq\;\int_{-\infty}^{-L+\lambda_{\max}^{-}(H_{1})T}\|\bm{w}(0,p)\|^{2}\mathrm{d}p.$
Integrating (5.5) in the time-space domain $(0,T)\times(-\infty,-L)$ gives
$\int_{-\infty}^{-L}\|\bm{w}(T,p)\|^{2}\mathrm{d}p-\int_{-\infty}^{-L}\|\bm{w}(0,p)\|^{2}\mathrm{d}p$
$=$ $-\int_{0}^{T}\int_{-\infty}^{-L}\partial_{p}\big(\bm{w}^{\dagger}H_{1}\bm{w}\big)\mathrm{d}p\mathrm{d}t=-\int_{0}^{T}\bm{w}^{\dagger}(t,-L)H_{1}\bm{w}(t,-L)\mathrm{d}t$ (5.6)
Thus, we have
$\left|\int_{0}^{T}\bm{w}^{\dagger}(t,-L)H_{1}\bm{w}(t,-L)\mathrm{d}t\right|\leq 2\int_{-\infty}^{-L+\lambda_{\max}^{-}(H_{1})T}\|\bm{w}(0,p)\|\mathrm{d}p.$ (5.7)
The proof for (5.4) is similar, which is omitted here. ∎
Lemma 5.3.
$\bm{w}:[0,T]\times\mathbb{R}\to\mathbb{C}^{N}$ be the solution to (2.4) and $\psi(p)$ satisfies (H1). Then it holds that
$\int_{0}^{T}\|\bm{w}(t,-L)\|^{2}\,\mathrm{d}t+\int_{0}^{T}\|\bm{w}(t,R)\|^{2}\,\mathrm{d}t\;\lesssim\;T\,\epsilon^{2}\,\|\bm{u}_{I}\|^{2},$ (5.8)
where the hidden constant is independent of $H_{1}$ and $\bm{u}_{I}$.
Proof.
Let $U(t)$ be the solution of the matrix ODE
$\partial_{t}U(t)=\mathrm{i}H_{2}(t)U(t),\qquad U(0)=I.$
Since $H_{2}(t)$ is Hermitian, $U(t)$ is unitary for all $t\in[0,T]$. Define
$\bm{v}(t,p):=U(t)^{-1}\bm{w}(t,p).$
Then $\|\bm{v}(t,p)\|=\|\bm{w}(t,p)\|$ and $\bm{v}$ solves
$\partial_{t}\bm{v}(t,p)=-\widetilde{H}_{1}\,\partial_{p}\bm{v}(t,p),\qquad\bm{v}(0,p)=\psi(p)\,\bm{u}_{I},$
where $\widetilde{H}_{1}:=U(t)^{-1}H_{1}U(t)$ is Hermitian and has the same eigenvalues as $H_{1}$. Since $H_{1}$ is time-independent, the spectrum of $\widetilde{H}_{1}$ is independent of $t$.
Diagonalizing $\widetilde{H}_{1}=Q\Lambda Q^{*}$ with a unitary $Q$ and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$, we set
$\bm{z}(t,p):=Q^{*}\bm{v}(t,p),\qquad\tilde{\bm{u}}_{I}:=Q^{*}\bm{u}_{I}.$
Then each component $z_{j}$ satisfies a scalar transport equation
$\partial_{t}z_{j}(t,p)=-\lambda_{j}\,\partial_{p}z_{j}(t,p),\qquad z_{j}(0,p)=\psi(p)\,\tilde{u}_{I,j},$
whose solution along characteristics is $z_{j}(t,p)=\psi(p+\lambda_{j}t)\,\tilde{u}_{I,j}.$
Evaluating at the boundaries p=Lp=-L and $p=R$ gives
$z_{j}(t,-L)=\psi(-L+\lambda_{j}t)\,\tilde{u}_{I,j},\quad z_{j}(t,R)=\psi(R+\lambda_{j}t)\,\tilde{u}_{I,j}.$
By the choice of $L,R$ and the definition of $\lambda_{\max}^{\pm}(H_{1})$, for every $j$ and $t\in[0,T]$ the characteristic footpoints satisfy
$L+\lambda_{j}t\in(-\infty,-L+\lambda_{\max}^{-}(H_{1})T],\quad R+\lambda_{j}t\in[R-\lambda_{\max}^{+}(H_{1})T,+\infty).$
Hence the decay condition (H1) yields
$|\psi(-L+\lambda_{j}t)|\leq 2\epsilon,\qquad|\psi(R+\lambda_{j}t)|\leq 2\epsilon,\quad 0\leq t\leq T,\ \forall j.$
Therefore
$|z_{j}(t,-L)|^{2}\leq 4\epsilon^{2}|\tilde{u}_{I,j}|^{2},\qquad|z_{j}(t,R)|^{2}\leq 4\epsilon^{2}|\tilde{u}_{I,j}|^{2}.$
Summing over $j$ and using the unitarity of $Q$ and $U(t)$, we obtain
$\|\bm{w}(t,-L)\|^{2}=\|\bm{z}(t,-L)\|^{2}=\sum_{j=1}^{N}|z_{j}(t,-L)|^{2}\leq 4\epsilon^{2}\sum_{j=1}^{N}|\tilde{u}_{I,j}|^{2}=4\epsilon^{2}\|\bm{u}_{I}\|^{2},$
and similarly $\|\bm{w}(t,R)\|^{2}\leq 4\epsilon^{2}\|\bm{u}_{I}\|^{2}$ for $0\leq t\leq T.$ Integrating in time yields (5.8). ∎
Similarly, the estimate for the periodic case follows an argument analogous to the previous proof; we therefore state it without proof in the following lemma.
Lemma 5.4.
Assume $\mathcal{W}$ is the solution to (5.2) with periodic boundary conditions. Then it holds
$\left|\int_{0}^{T}\mathcal{W}^{\dagger}(t,-L)H_{1}\mathcal{W}(t,-L)\mathrm{d}t\right|\lesssim\,\int_{-L}^{-L+\lambda_{\max}^{-}(H_{1})T}\|\mathcal{W}(0,p)\|^{2}\mathrm{d}p+\int_{R-\lambda^{+}_{\max}(H_{1})T}^{R}\|\mathcal{W}(0,p)\|^{2}\mathrm{d}p.$ (5.9)
Lemma 5.5.
Assume $\mathcal{W}$ is the solution to (5.2) with periodic boundary conditions. Then it holds
$\int_{0}^{T}\|\mathcal{W}(t,-L)\|^{2}\,\mathrm{d}t+\int_{0}^{T}\|\mathcal{W}(t,R)\|^{2}\,\mathrm{d}t\;\lesssim\;T\,\epsilon^{2}\,\|\bm{u}_{I}\|^{2}.$ (5.10)
Then, we get the error between $\bm{w}(t,p)$ and $\mathcal{W}(t,p)$.
Theorem 5.3.
Let $\mathcal{E}(t,p)=\bm{w}(t,p)-\mathcal{W}(t,p)$. Assume $\psi$ satisfies (H1). It follows that
$\|\bm{w}(T,p)-\mathcal{W}(T,p)\|_{L^{2}((-L,R))}\lesssim\epsilon\|\bm{u}_{I}\|.$
Proof.
The error function $\mathcal{E}(t,p)$ satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}=-H_{1}\partial_{p}\mathcal{E}+\mathrm{i}H_{2}\mathcal{E},\quad\mathcal{E}(0,p)=\bm{0}.$ (5.11)
As in (5.1), by testing (5.11) against $\mathcal{E}^{\dagger}$ and integrating by parts, we arrive at
$\|\mathcal{E}(T,p)\|_{L^{2}((-L,R))}^{2}\lesssim$ $\left|\int_{0}^{T}\bm{w}^{\dagger}(t,-L)H_{1}\bm{w}(t,-L)\mathrm{d}t\right|+\left|\int_{0}^{T}\bm{w}^{\dagger}(t,R)H_{1}\bm{w}(t,R)\mathrm{d}t\right|$
$+\left|\int_{0}^{T}\mathcal{W}^{\dagger}(t,L)H_{1}\mathcal{W}(t,L)\mathrm{d}t\right|+\frac{1}{2T}\int_{0}^{T}\|\bm{w}(t,-L)\|^{2}\,\mathrm{d}t$
$+\frac{1}{2T}\int_{0}^{T}\|\bm{w}(t,R)\|^{2}\,\mathrm{d}t+\frac{1}{2T}\int_{0}^{T}\|\mathcal{W}(t,-L)\|^{2}\,\mathrm{d}t+\frac{1}{2T}\int_{0}^{T}\|\mathcal{W}(t,R)\|^{2}\,\mathrm{d}t.$
which completes the proof by applying Lemma 5.25.5. ∎
5.2.2 Main proof of Theorem 2.1
For brevity, we set a=La=-L and $b=R$. Define the complex $N_{p}$-dimensional space
$X_{N_{p}}:=\text{span}\big\{\phi_{l}(p)=\mathrm{e}^{\mathrm{i}\mu_{l}(p-a)}:0\leq l<N_{p}\big\},$
where $\mu_{l}=\frac{2\pi}{b-a}(l-\frac{N_{p}}{2})$. Let $\bm{w}(t,p)=[w_{1}(t,p),\cdots,w_{2N}(t,p)]^{\top}$ be the solution to (2.4). The approximate solution $\bm{w}_{h}\in(X_{N_{p}})^{2N}$ is then given in (2.26).
Define the $L^{2}$-orthogonal projection $P_{h}:L^{2}((a,b))\to X_{N_{p}}$ , given by
$P_{h}u=\sum_{k=0}^{N_{p}-1}\hat{u}_{l}\phi_{l}(p),\qquad\hat{u}_{l}=\frac{1}{b-a}\int_{a}^{b}u\mathrm{e}^{-\mathrm{i}\mu_{l}(p-a)}\mathrm{d}p.$
Next, we consider the Fourier interpolation denoted by $\Pi$, that is
$\Pi u(p)=\sum_{l=0}^{N_{p}}\tilde{u}_{l}\phi_{l}(p),\quad\tilde{u}_{l}=\frac{1}{N_{p}}\sum_{k=0}^{N_{p}}u(p_{k})\mathrm{e}^{-\mathrm{i}\mu_{k}(p_{l}-a)},\quad 0\leq l<N_{p}.$
Then, one has the following estimates for $u\in H^{r}_{p}((a,b))$ [Shenspectral]
$\|P_{h}u-\Pi u\|_{L^{2}((a,b))}\lesssim(\triangle p)^{r}\|u^{(r)}\|_{L^{2}((a,b))}.$ (5.12)
Using the triangle inequality, one has
$\|\bm{w}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}\leq\|\bm{w}(\cdot,p)-\mathcal{W}(\cdot,p)\|_{L^{2}((a,b))}+\|\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}.$ (5.13)
According to Theorem 5.3, it is sufficient to prove the second term. By the triangle inequality, the second part of (5.13) can be split as
$\|\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}\leq\|\mathcal{W}(\cdot,p)-P_{h}\mathcal{W}(\cdot,p)\|_{L^{2}((a,b))}+\|P_{h}\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}.$
For the first term on the right-hand side, the standard projection error estimate for the discrete Fourier transform yields
$\|\mathcal{W}-P_{h}\mathcal{W}(\cdot,p)\|_{L^{2}((a,b))}\lesssim\Big(\frac{b-a}{N_{p}}\Big)^{k}\|\mathcal{W}^{(k)}(\cdot,p)\|_{L^{2}((a,b))}\lesssim\Big(\frac{b-a}{N_{p}}\Big)^{k}\|\psi^{(k)}\|_{L^{2}((a,b))}\|\bm{u}_{I}\|.$
The second inequality follows by expanding $\mathcal{W}$ and $\psi$ in the periodic Fourier basis and observing that, for each Fourier mode, the evolution operator $e^{\,\mathrm{i}(-\xi H_{1}+H_{2})t}$ is unitary, so that $\|\partial_{p}^{k}\mathcal{W}(t,\cdot)\|_{L^{2}((a,b))}$ is controlled by $\|\psi^{(k)}\|_{L^{2}((a,b))}\,\|\mathbf{u}_{I}\|$ with a constant independent of $t$. For the second term, by definition, one gets
$P_{h}\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)=\sum\limits_{l=0}^{N_{p}-1}(\hat{\bm{w}}_{l}(t)-\tilde{\bm{w}}_{l,h}(t))\phi_{l}(p),$
where $\hat{\bm{w}}_{l}=\int_{a}^{b}\bm{w}e^{-i\mu_{l}(p-a)}\mathrm{d}p/(b-a)$, which gives
$\|P_{h}\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}^{2}=\sum\limits_{l=0}^{N_{p}-1}\|\hat{\bm{w}}_{l}(t)-\tilde{\bm{w}}_{l,h}(t)\|^{2}.$
Let $\hat{\bm{e}}_{l}=\hat{\bm{w}}_{l}-\tilde{\bm{w}}_{l,h}$. It is easy to check that
$\begin{cases}\dfrac{\mathrm{d}}{\mathrm{d}t}\hat{\bm{e}}_{l}(t)=-\mathrm{i}(\mu_{l}H_{1}-H_{2})\hat{\bm{e}}_{l}(t)\\ \hat{\bm{e}}_{l}(0)=(\hat{\psi}_{l}-\tilde{\psi}_{l})\bm{u}_{I}\end{cases}\qquad l=0,\cdots,N_{p}-1.$
Therefore, we have
$\|P_{h}\mathcal{W}(\cdot,p)-\bm{w}_{h}(\cdot,p)\|_{L^{2}((a,b))}^{2}$ $=\sum_{l=0}^{N_{p}-1}\|\hat{e}_{l}\|^{2}=\sum_{l=0}^{N_{p}-1}|\hat{\psi}_{l}-\tilde{\psi}_{l}|^{2}\|\bm{u}_{0}\|^{2}$
$=\|P_{h}\psi-\Pi\psi\|^{2}_{L^{2}((a,b))}\|\bm{u}_{I}\|^{2}.$ (5.14)
The proof is finished by using (5.12) and (2.24).
6 Discussion
The LCHS method proposed in [ALL2023LCH] is closely related to the Schrödingerization framework. For example, by taking $\psi(p)=\mathrm{e}^{-|p|}$ in (2.4), one applies the continuous Fourier transform on $p$ to get the Schrödinger-type system
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\bm{w}}=\mathrm{i}(\xi H_{1}+H_{2})\hat{\bm{w}},\qquad\hat{\bm{w}}(0)=\frac{1}{\pi(1+\xi^{2})}\bm{u}_{I}.$
This, together with the recovery formula, yields
$\bm{u}=\bm{w}(t,0)=\int_{\mathbb{R}}\frac{1}{\pi(1+\xi^{2})}\mathcal{T}\exp\!\Big(\mathrm{i}\int_{0}^{t}\xi H_{1}(s)+H_{2}(s)\,\mathrm{d}s\Big)\bm{u}_{I}\,\mathrm{d}\xi,$
when the eigenvalues of $H_{1}$ are non-positive, where $\mathcal{T}\exp$ is the time-ordering exponential operator, which is consistent with the exact representation in [ALL2023LCH, Theorem 1]. Therefore, the two approaches share the same foundation at the analytical level.
Due to the same foundation, the LCHS in [ALL2023LCH] is also a first-order method. Subsequent improvements on the precision are presented in [ACL2023LCH2], with their relation to our smooth extension covered in Section 3.3. The differences between the Schrödingerization and the LCHS arise in the discretization strategies applied to the auxiliary variable $p$. In particular, Schrödingerization leads to the matrix-query complexity of order
$\mathcal{O}\!\big(\log(1/\varepsilon)\big),$
whereas the optimized LCHS method in [ACL2023LCH2] achieves
$\mathcal{O}\Big((\log(1/\varepsilon))^{1/\beta}\Big),\qquad\beta\in(0,1).$
in the time-dependent case. Clearly, the latter one has sub-optimal dependence on matrix queries since $\beta$ cannot be exactly equal to 1.
During the revision stage of this work, we were informed of a recent study [Low25LCHS], which introduces a generalized approximate version of LCHS. This extension incorporates a kernel function with exponential decay – which inspired our own work – also allowing for a quantum ODE algorithm that can achieve optimal precision dependence.
Acknowledgements
We kindly thank Dr. Guang Hao Low for pointing out to us some errors in the earlier version of this paper, where our earlier result gave a scaling that was near-optimal instead of optimal. This helped us to provide an improved result and to achieve optimality.
SJ and NL are supported by NSFC grant No. 12341104, the Shanghai Jiao Tong University 2030 Initiative and the Fundamental Research Funds for the Central Universities. SJ was also partially supported by the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102). NL also acknowledges funding from NSFC grant No.12471411 and the Science and Technology Program of Shanghai, China (21JC1402900). CM was partially supported by NSFC grant No. 12501607, the Science and Technology Commission of Shanghai Municipality (No.22DZ2229014), the China Postdoctoral Science Foundation (No. 2023M732248) and the Postdoctoral Innovative Talents Support Program (No. BX20230219). YP was supported by the Xiangtan University 2025 Postgraduate Innovation Project (No. XDCX2025Y216). YY was supported by NSFC grant No. 12301561, the Key Project of Scientific Research Project of Hunan Provincial Department of Education (No. 24A0100), the Science and Technology Innovation Program of Hunan Province (No. 2025RC3150) and the 111 Project (No. D23017).